### 20 Jan 2024 v1

# StickerConv: Generating Multimodal Empathetic Responses from Scratch

Abstract
Stickers, while widely recognized for enhancing empathetic communication in online interactions, remain underexplored in current empathetic dialogue research. In this paper, we introduce the Agent for StickerConv (Agent4SC), which uses collaborative agent interactions to realistically simulate human behavior with sticker usage, thereby enhancing multimodal empathetic communication. Building on this foundation, we develop a multimodal empathetic dialogue dataset, StickerConv, which includes 12.9K dialogue sessions, 5.8K unique stickers, and 2K diverse conversational scenarios, specifically designs to augment the generation of empathetic responses in a multimodal context. To leverage the richness of this dataset, we propose PErceive and Generate Stickers (PEGS), a multimodal empathetic response generation model, complemented by a comprehensive set of empathy evaluation metrics based on LLM. Our experiments demonstrate PEGS’s effectiveness in generating contextually relevant and emotionally resonant multimodal empathetic responses, contributing to the advancement of more nuanced and engaging empathetic dialogue systems. Our project page is available at https://neu-datamining.github.io/StickerConv.
StickerConv: Generating Multimodal Empathetic Responses from Scratch
[…]
footnotetext: Corresponding author.**footnotetext: Equal contribution.
1 Introduction
Increasing research indicates that utilizing stickers in online chats can effectively alleviate stress, augment personal happiness, and notably boost empathy Kariko and Anasih (2019); Akram and Drabble (2022); Jiang et al. (2023). Prior studies on stickers primarily concentrated on sentiment analysis Ge et al. (2022) and recommendation systems Gao et al. (2020); Zhang et al. (2022), overlooking their vast potential in empathetic response generation. Most empathetic response generation tasks focus solely on textual modality Fu et al. (2023); Lee et al. (2022); Jiang et al. (2023), yet stickers in chats convey more abundant and intuitive emotional information, enhancing the expressiveness and emotional depth of responses.
Figure 1: An example of multimodal conversation in the StickerConv. Both parties can utilize the stickers to express their emotions, which enhances interactivity and expression. Assistant can empathize with the user according to the conversation (green text).
Acknowledging stickers as a crucial non-textual modality and incorporating them into empathetic response generation offers novel perspectives and methodologies for empathetic dialogue tasks Zech and Rimé (2005); Rashkin et al. (2019). Specifically, integrating text and stickers paves the way for a new multimodal empathetic communication approach, targeting the generation of more diverse and high-quality empathetic responses.
A primary challenge in integrating stickers into empathetic response generation is developing a high-quality dataset to support this innovative multimodal communication. Traditional dataset construction methods, including expensive human crowdsourcing, privacy issues Xu et al. (2022), the difficulty in acquiring data from social software companies, and the variable quality of Internet-crawled data Foerderer (2023), face significant limitations.
To address this, leveraging large language models (LLMs) for dataset construction is proposed. LLMs, with their extensive world knowledge and text processing capabilities (OpenAI, 2023; Touvron et al., 2023a, b), demonstrate near-human annotation abilities Wang et al. (2023d); Pangakis et al. (2023). However, LLMs have limitations in empathetic tasks, excelling in responding to explicit human instructions but lacking in proactivity Wu et al. (2023a); Yin et al. (2023), a critical aspect of empathy. Empathy necessitates understanding others’ emotions and the ability to actively express support and understanding Zech and Rimé (2005); Sharma et al. (2020a). To mitigate this, we introduce a multi-agent system based on LLMs, Agent for StickerConv (Agent4SC). This system, through inter-agent interactions, utilizes stickers to simulate human-like dialogue scenarios. It not only generates text responses but also strategically selects suitable stickers, thereby effectively enhancing empathy.
Based on the superior capabilities and interactions of Agent4SC, we built a multimodal empathetic dataset, StickerConv, that comprise 12.9K dialogue sessions and 5.8K unique stickers. StickerConv boasts an average of 5.22 stickers per dialogue session. Figure 1 depicts an example of conversations in our dataset. To the best of out knowledge, this is the first multimodal empathetic dialogue dataset, with the particular utility of sticker as non-textual modal information to better facilitate empathy.
Although Agent4SC exhibits sound multimodal empathetic dialog response generation capability, it is limited by expensive inference costs, while the retrieval approach leads to poor performance in new scenarios. To further advance the research on multimodal empathetic dialog, based on StickerConv, we develop a end-to-end large multimodal dialog model, PEGS, with the ability to PErceive and Generate Stickers, and has a significantly lower deployment and inference cost. Beyond the general ability to generate textual empathetic responses, PEGS receives multimodal inputs and autonomously generates stickers based on the emotional and contextual aspects of the dialogue at the appropriate moment. Furthermore, to simulate human communications on social media in the real world, our model supports interleaved multiple image and text inputs.
Due to potential misalignments between the modality quantities of predicted responses and those of the golden responses, reliance on a singular modal metric may lead to significant distortions in evaluation results. Empathy, being inherently subjective, presents challenges in its direct quantification. In this context, we introduce a novel multimodal empathetic response evaluation method. This method enables the integration of multimodal information, emphasizing three critical dimensions: empathy, consistency, and preference ranking. Leveraging the world knowledge and anthropomorphic capabilities of LLMs, our method offers robust support for evaluating multimodal empathetic responses.
In conclusion, the main contributions of this work are as follows:
  • We introduce Agent for StickerConv (Agent4SC), a LLM-based multi-agent system, integrates stickers intelligently into empatheric dialogues. Agent4SC ensures the stickers are contextually consistent, varied, and empathetically aligned with human interaction. Leveraging the Agent4SC, we built a multimodal empatheric dialogue dataset, StickerConv, comprising 12.9K dialogue sessions and 5.8K unique stickers.
  • We design PErceive and Generate Stickers (PEGS), an empathetic dialogue response generation model that, within its end-to-end framework, intuitively incorporates stickers based on the emotional and contextual dynamics of a dialogue. PEGS adeptly processes multimodal inputs, generating empathetic textual responses and appropriately using stickers to enrich these replies.
  • We propose a multimodal empathetic response evaluation approach. This method integrates multimodal information and concentrates on three principal dimensions: empathy, consistency, and preference ranking.
2 Related Work
2.1 Empathetic Response Generation
The purpose of empathetic response generation is to enable machines to understand and respond to human emotions and feelings. In this field, The creation of the EMPATHETICDIALOGUES dataset Rashkin et al. (2019) laid the foundation. Subsequently, the empathetic listener mixture model Lin et al. (2019), which emphasizes understanding and appropriately responding to different emotions, and studies focusing on emotion imitation and emotion cause words to enhance the effectiveness of empathetic dialogue Majumder et al. (2020) have achieved notable results. Additionally, the introduction of the EPITOME conceptual framework Sharma et al. (2020b), a classification system for empathetic response intents Welivita and Pu (2020), research on generating empathetic responses based on commensense knowledge Sabour et al. (2021), and the proposal of the multifactorial hierarchical framework CoMAE Zheng et al. (2021) have provided new perspectives and methods for the field.
Despite these advancements, the research in this domain is not without its limitations. Primarily, the field has been domainated by small-scale models, with limited application of large language models, such as ChatGPT OpenAI (2023), in the generation of empathetic responses Lee et al. (2022); Zhou et al. (2023b). Secondly, the focus has predominantly been on text-based empathetic responses, thus neglecting the rich potential of multimodal information, such as image. Emerging studies underscore the significance of such multimodal data, especially stickers, in alleviating stress, enhancing well-being, and fostering emotional resonance and social connectivity Kariko and Anasih (2019); Akram and Drabble (2022); Jiang et al. (2023). Finally, the subjectivity and multidimensionality of empathetic tasks make it challenging to quantitatively assess their effectiveness Fu et al. (2023); Lee et al. (2022), especially in the case of multimodal empathetic responses, posing a significant challenge in the field of empathetic response generation.
2.2 Large Multimodal Models
Multimodal Perception
Large language models (LLMs), such as ChatGPT (OpenAI, 2023), LLaMA (Touvron et al., 2023a, b), demonstrate powerful capabilities in dialog interaction and instruction following, and recent researches have extended LLMs to multimodal domains. Flamingo (Alayrac et al., 2022) exhibits promising zero-shot and few-shot multimodal understanding capability by adding a cross-attention layer to connect the frozen vision encoder with the LLM. BLIP (Li et al., 2023b; Dai et al., 2023), MiniGPT-4 (Zhu et al., 2023; Chen et al., 2023) and LLaVA (Liu et al., 2023b, a) bridge the frozen vision encoder and the LLM through a small intermediate model.
Multimodal Generation
The birth and development of the Diffusion Models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Rombach et al., 2022) significantly contribute to the evolution of image generation. Koh et al. (2023) for the first time explore the mapping of the output of LLMs into the input space of the vision decoder, empowering LLMs with image generation capability. GILL (Koh et al., 2023), NExT-GPT (Wu et al., 2023b), MiniGPT-5 (Zheng et al., 2023) align LLMs to frozen vision decoders through an encoder-decoder transformer, while SEED (Ge et al., 2023) employs a learnable Q-Former. In contrast to previous works which favor realistic images, our target is to generate stickers, which are abstract and exhibit distinct emotional tendencies.
2.3 LLM-Based Agents
In the realm of artificial intelligence, Large Language Model (LLM)-based agents epitomize a cutting-edge paradigm, leveraging the robust computational prowess of LLMs. These agents are not merely confined to problem-solving and task execution; their capabilities extend to intricate reasoning and sophisticated interactions, as elucidated in recent studies Wang et al. (2023b); Li et al. (2023a). The versatility of LLM agents is further demonstrated through their adaptability in diverse domains such as software development Qian et al. (2023) and scientific research Boiko et al. (2023).
A remarkable aspect of LLM-based agents is their ability to emulate complex human social behaviors. This includes a spectrum of interactions ranging from social engagements Park et al. (2023); Tu et al. (2023), to simulating multifaceted scenarios like elections Argyle et al. (2022), debates Wang et al. (2023a), and consumer activities Wang et al. (2023c), Such capabilities underscore the potential of LLM agents in understanding and replicating human social dynamics.
The most transformative application of LLM-based agents, however, lies in their utility for generating training data for diverse AI models. Studies Wang et al. (2023d); Peng et al. (2023); Zhou et al. (2023a) have been instrumental in showcasing the efficacy of these agents in producing high-quality data, paralleling human annotators, yet at a substantially lower cost. This breakthrough has redefined the paradigm for creating training datasets, offering a more practical and cost-effective solution in the field of AI model development.
Figure 2: The Overview of Agent for StickerConv.
3 Agent for StickerConv
The limitations of Large Language Models (LLMs) and Language Model Multimodalities (LMMs), are twofold: (1) They struggle with capturing nuanced human emotions and providing empathetic responses. (2) While adept at responding to explicit human instructions, they lack the initiative to act beyond these directives.
This inherent disadvantage hinders their performance in tasks that require empathy, making them less effective in this area compared to others. To bridge this gap, we propose the Agent for StickerConv (Agent4SC), a novel LLM-based multi-agent. This system is designed to simulate human dialogue by leveraging the collaborative efforts of multiple modules, along with the use of stickers, to produce emotional and diverse empathetic responses.
The architecture of Agent4SC is depicted in Figure 2 and comprises two key agents. The User Agent, forming the core of the system, is an intricate assembly of several modules. It replicates the human processes of observation and reflection through the interplay between its Memory and Plan modules. This design effectively addresses the LLMs’ shortcomings in comprehending subtle human emotions. The insights gained from observation and reflection are articulated via the Action module. Consequently, the Agent is more adept at generating human-like emotional responses. The entire process is orchestrated by the Profile module, which imparts distinctiveness in reflection and behavior to different User Agents.
Furthermore, Agent4SC adopts a pioneering approach by treating stickers as a Tool of communication. Agents autonomously select appropriate stickers for use, mirroring human behavior in conversations. This function is facilitated through the integrated process of observation, reflection, and action. The User Agent ensures consistent performance and quality control with the support of the Manager Agent. Agent4SC is not just a static tool but a dynamic participant in conversations, capable of grasping the subtleties of human emotions and the appropriate moments for empathetic interaction.
3.1 Profile Module
The profile module serves as the foundation of the Agent4SC, shaping the personality and behavioral patterns of users in empathetic interactions. It comprises two components: Persona and Situation. The Persona section vividly portrays the user’s character traits, life background, and experiences, whereas the Situation part focuses on the user’s current circumstances, emotional state, and potential challenges they may encounter.
To develop rich and diversified user profiles, we initially manually crafted profiles with a variety of emotioinal distributions, subsequently expanding them to 2K unique profiles through the SELF-INSTRUCT Wang et al. (2023d) method by ChatGPT-3.5-turbo. Our constructed profile archive features a comprehensive emotional distribution, as illustrated in Figure 5. We believe that emotionally rich and content-diverse user profiles will enable the agent system to more authentically simulate human responses in empathetic dialogues.
3.2 Tool Module
To effectively simulate human behavior in using stickers for communication, we developed a Tool Module, converting the SER30K dataset into a functional tool. This dataset comprises 1,887 themes and 30,739 emotion-labeled stickers (Liu et al., 2022), such as happiness, anger, neutral, disgust, fear, sadness and surprise. In integrating it with the agent system, we encountered several challenges: (1) The dataset contains a significant number of stickers unsuitable for casual conversation or empathetic communication. (2) Stickers are annotated with a single emotional tag from a set of seven, lacking comprehensive content descriptions and a detailed emotional analysis. This limitation hinders the agent’s ability to appropriately utilize stickers based solely on these tags. (3) Humans selecting stickers in conversation typically do not choose from a database containing tens of thousands of sticker. Instead, they prefer to pick from a customized and smaller-scale sticker library.
To overcome these challenges, we design a series of sophisticated sticker processes to refashion sticker into an effective tool for accurate agent application:
Filter
Leveraging the LLaVa-v1.5-13b (Liu et al., 2023a), we crafted a nuanced three-step Chain of Thought Wei et al. (2023) procedure to meticulously scrutinize each sticker in SER30K, as shown in Figure 9. This approach entailed the model initially summarizing the sticker’s content, subsequently assessing its relevance to casual conversation, and finally determining suitability through iterative responses. This strategy not only affirmed selection accuracy but also mitigated model-induced randomness and hallucination discrepancies.
Annotation
For filtered stickers, a second iteration of the CoT process was executed using the same model. This phase involved responding to five pivotal queries, covering aspects like the sticker’s content, emotional conveyance, applicable scenarios, emotional tone, and potential satirical or humorous connotations (shown in Figure 9). These responses yielded comprehensive insights into each sticker, subsequently formatted into Q&A pairs, awaiting further processing.
Knowledge
Post-acquisition of intricate sticker details, the ChatGPT-3.5-Turbo model was engaged for an additional processing phase. Relying on the established Q&A pairs, this model adeptly deduced each sticker’s specific contextual suitability, culminating in the distillation of three fundamental knowledge: description, emotion, recommendation. These synthesized insights are instrumental in enriching agent’s understanding of sticker, and enable them to apply stickers correctly.
Split
In this phase of refining the SER30K sticker dataset for integration into our agent system, we focus on dividing the database into smaller, more manageable subsets. Each subset is transformed into a vector database, where the embeddings represent a unique blend of the original emotional labels of the stickers and the enriched emotional knowledge extracted in the previous steps.
After the above-mentioned processes, the original SER30K dataset has been transformed into a series of refined, smaller-scale vector databases. However, a crucial step is still required to make it a directly usable tool for agents. To address this, we have defined a specific module operation, termed "Retrieve". In this process, when an agent intends to use a sticker, it queries its accessible sub-vector database with its intended use and desired emotional expression of the sticker as search criteria. The system then retrieves and returns the top-K stickers with the highest similarity.
3.3 Plan Module
The Plan module is a core component of the Agent4SC, responsible for orchestrating the behavior of agents in empathetic interactions. To ensure that agents mimic human behavior in using stickers, we have specially designed three key actions to ensure the timely use of stickers, their coherence with the context, and their effective role in empathy.
  • Intention: Under the control of Profile module, the agent assesses the potential for using emotive stickers based on interaction history and observational data. The prompt template for this action is shown in Figure 11.
  • Query: If the agent intends to use a sticker, it will describe the emotional content of the sticker in detail. This description will be influenced by the Profile module, observation, and the specific intent behind using the sticker. The prompt template for this action is shown in Figure 12.
  • Select: The agent sends the query generated in the previous step to the sticker database accessible to it, retrieving the top-K stickers with the highest similarity. Subsequently, the agent combines the candidated sticker’s index, description, emotion and recommendation into a list. From this list, the agent selects the sticker it wishes to use. The prompt template for this action is shown in Figure 13.
3.4 Memory Module
In empathetic interactions, the coherence of communication and the consistency of personality traits are fundamental to the quality of responses. This necessitates that agents not only learn and provide feedback through observation during the interaction process, but also maintain long-term consistency in their personality traits. We adopted a strategy similar to that of Generative Agent Park et al. (2023): for short-term memory, we update the agent’s core traits based on the profile and observations; for long-term memory, the agent combines continuous interaction with external entities and regular self-reflection to maintain consistency.
3.5 Action Module
The Action module serves as the window and bridge for communication with the external world. It is responsible for receiving instructions and inputs from other parts of the system and executes two main actions based on this information:
  • Message: generates text responses based on the Agent’s Profile and Memory module.
  • Sticker: engages in sticker usage as dictated by the Plan module.
3.6 Manager Agent
The Manager Agent, a critical componet in enhancing the performance of the Agent4SC, plays a significant role in ensuring interaction quality and dialogue consistency. This agent comprises two main modules:
Quality Assurance
This module is crucial for upholding high dialogue standards, It accomplishes this by: (1) Evaluating the relevance of stickers within the conversation, ensuring that they contribute meaningfully to the dialogue. (2) Regulating the use of stickers to prevent their overuse, thus maintaining the integrity and flow of the conversation. (3) Performing regular checks on dialogue quality, with a special focus on the appropriateness of response length and the avoidance of redundant information.
Process Control
This module takes charge of guiding the dialogue’s progression. It leverages historical interactions as a reference to make informed decisions on the optimal timing for concluding dialogues.
Figure 3: The chart of emotional distribution in the choice of stickers between users and the system. This chart revealed a striking trend: users have a significant preference for stickers that convey negative emotions, in contrast to the system. The system predominantly utilizes stickers expressing neutral and positive emotions. This comparison not only reflects the distinct emotional expression preferences between users and the system but also highlights the system’s active and supportive role in interactions.
3.7 The StickerConv Dataset
The Agent4SC facilitates empathetic interaction with humans through the collaborative interplay of its modules, and possesses the capability to utilize stickers. Leveraging this, we propose the construction of a large-scale multimodal empathetic dialogue dataset, StickerConv, with stickers serving as a non-textual modality.
Figure 4: The overview architecture of PEGS. The input images are jointly encoded by image encoder, Q-Former and a linear layer. Vicuna is used as the language model. To empower the language model to generate images, the output of the LLM is first mapped into the input space of the image decoder through a feature mapper, and then the frozen image decoder is employed to generate images.
In the dataset creation process, two roles are designated: User and System. The User, acting as the dialogue initiator, is assigned a profile randomly selected from 2,000 personality profiles generated as described in Section 3.1. The System, in the roles of a listener and empathizer, is assigned a predefined positive profile. Regarding the sticker tool, we have processed and subdivided the sticker dataset into 100 smaller vector databases. In each dialogue session, both User and System have access to one randomly assigned vector database from this pool. This arrangement is designed to better simulate human sticker usage behavior while enhancing the diversity and richness of stickers. It is important to note that vector databases numbered 1 to 80 are exclusively employed for the creation of training and cross-validation sets, whereas those numbered 81 to 100 are reserved for the construction of the test set.
StickerConv consists of 12,931 dialogue sessions, 67,505 stickers (unique 5.8K stickers) and over 2K conversation topic. The detailed statistical information of our dataset is shown in Table 4. On average, each dialogue session in StickerConv has 5.22 stickers and 5.49 turns. The emotional label distribution of the stickers used by User and System is shown in Figure 3. This analysis reveals notable differences in the emotional distribution of sticker usage, attributable to the distinct roles of the User and the System.
4 PEGS
We devise an empathetic large multimodal dialog model, PEGS, with the ability to PErceive and Generate Stickers. Figure 4 illustrates the overall structure of our model. Technically, we utilize ViT-G/14 from EVA-CLIP (Fang et al., 2023), Q-Former from BLIP-2 $\text{OPT}_{\text{2.7B}}$ (Li et al., 2023b), and a linear layer to encode images. Vicuna (Chiang et al., 2023), a widely used language model in LMMs, is employed for language modeling. We adopt Stable Diffusion (SD) (Rombach et al., 2022) as our image decoder.
4.1 Multimodal Input Perception
Multimodal Encoding
With reference to existing works (Li et al., 2023b; Dai et al., 2023; Zhu et al., 2023), we convert the multimodal inputs into feature vectors that can be solved by LLM. Specially, each text token is embedded into a vector $e_{t}\in\mathbb{R}^{1\times d}$, while each image is first encoded by a pre-trained vision encoder, and then a aligned feature vector $e_{v}\in\mathbb{R}^{32\times d}$ is obtained via Q-Former and a linear projection layer.
Training
Our effort is built on BLIP-2 (Li et al., 2023b), which has captured vision-language knowledge from numerous aligned image-text pairs. Zhu et al. (2023) have proven the effectiveness that freezing the pre-trained vision encoder, Q-Former, and LLM, while only pre-training a linear projection layer. As shown in Figure 4(a), we train the input linear projection layer using LAION115M, a sub-dataset of LAION400M (Schuhmann et al., 2021). The pre-training process undergoes 80,000 steps with a batch size of 32.
4.2 Multimodal Output Generation
Expanding Vocabulary
Given original LLM vocabulary $V$ of size $|V|$, to enable the LLM to determine the timing of generating visual responses and to obtain feature vectors for image generation, we extend the vocabulary with an additional tokens set $V_{\mathrm{img}}=\{\mathrm{[IMG1]},\mathrm{[IMG2]},\ldots,\mathrm{[IMG\{}r\mathrm{\}]}\}$. Therefore, the extended vocabulary $V^{*}=V\cup V_{\mathrm{img}}$. We denote the original word embedding matrix as $E\in\mathbb{R}^{|V|\times d}$. For the embeddings matrix $E^{*}\in\mathbb{R}^{|V^{*}|\times d}$ of the extended vocabulary $V^{*}$, the embeddings $E_{\mathrm{img}}\in\mathbb{R}^{r\times d}$ of the added special token $V_{\mathrm{img}}$ are randomly initialized and the embeddings $E$ of the original token are preserved:
$\centering E^{*}[0:|V|,:]=E\@add@centering$
Text Generation
Receiving multimodal inputs, the target is to generate joint sequences of text tokens and visual tokens $\{\mathrm{[IMG\{}i{\}]}\}_{i=1}^{r}$. Specifically, the generated token can be represented as $U=\{u_{1},\ldots,u_{k}\}$, where $u_{i}\in V^{*}$. The loss function $\mathcal{L}_{\mathrm{lm}}$ is defined as:
$\centering\mathcal{L}_{\mathrm{lm}}=-\sum_{i=1}^{k}\log p(u_{i}|s,u_{1},\ldots,u_{i-1};\theta,E_{\mathrm{img}})\@add@centering$
where $s=\{e_{m}^{(1)},e_{m}^{(2)},\ldots,e_{m}^{(l)}\}$ and $m\in\{t,v\}$ denoting the modality.The original LLM weights $\theta$ are kept frozen, and we only update $E_{\mathrm{img}}$.
Image Generation
Inspired by GILL (Koh et al., 2023), we align the hidden states $h_{\mathrm{gen}}$ corresponding to the output visual tokens $\{\mathrm{[IMG\{}i{\}]}\}_{i=1}^{r}$ into the input space of the image decoder. Specifically, we connect them through a feature mapper module, containing two linear layers and a 4-layer encoder-decoder transformer model with a learnable queries feature $q$. For the given image caption $c$ (and its emotion $e$ if available), our target is to minimize the MSE loss between their embeddings derived from the frozen pre-trained SD text encoder $\tau$ and the projected representations:
$\centering\mathcal{L}_{\mathrm{gen}}=\lVert\theta_{\mathrm{Mapper}}(h_{\mathrm{gen}},q)-\tau(c,[e])\rVert_{2}^{2}\@add@centering$
Training
Distinguished from realistic images, stickers have abstract nature and rich emotional expressions. Generic text-to-image models perform well in various scenarios, but they are flawed in sticker generation. Constrained by limited sticker data, there are not enough image-text pairs to turn the base text-to-image model to the style of stickers, and to pre-train for multimodal generation. As a remedy, we adopt a variant of CuteYukiMix [^1] https://civitai.com/models/28169/cuteyukimixadorable-style, a SD model trained on cute cartoon data, as the pre-trained text-to-image model. In the pre-training phase, our objective is to align the representations with the text embeddings (Figure 4(b)), so we train our model on the captions from LAION115M. The pre-training process undergoes 50,000 steps with a batch size of 32.
However, pre-trained SDs (focusing on obvious features, such as characters’ bodies, costumes) are insensitive to sentiment conditions, and it is difficult to directly perform end-to-end instruction tuning to learn emotion elements, we inserted a supervise fine-tuning (SFT) procedure on the sticker-text pairs with emotion to efficiently cause them to focus on emotion elements.
4.3 Joint Learning
Pre-trained LLMs have powerful text interaction capability, but perform poorly on open-domain conversations. We further utilize the constructed StickerConv for jointly fine-tuning of the entire model to achieve the capabilities of open-domain empathetic multimodal perception and generation. The instruction prompts we used are shown in Appendix C.4. As shown in Figure 4(c), we train our model in an end-to-end manner, using LoRA (Hu et al., 2021) to synchronize the update of a limited number of parameters in the LLM with the input linear projection layer and the feature mapper. The overall loss function $\mathcal{L}$ can be represented as:
$\centering\mathcal{L}=\lambda_{1}\times\mathcal{L}_{\mathrm{lm}}+\lambda_{2}\times\mathcal{L}_{\mathrm{gen}}\@add@centering$
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters. For $\mathcal{L}_{\mathrm{gen}}$, one of its inputs — emotion $e$ is available.
5 Experiments
Experiments are conducted on StickerConv. We perform response predictions for all turns of each dialogue and consider all previous turns as context.
Models IP EX ER diff-IP diff-EX diff-ER dist-1 dist-2 dist-3
Vicuna-7B 0.1857 0.2405 0.9509 0.3681 0.9745 0.2651 0.7995 0.9783 0.9936
ChatGLM3-6B 0.1632 0.2699 0.9768 0.344 0.9993 0.2629 0.7854 0.9776 0.9939
PEGS 0.1723 0.4458 0.9724 0.3113 0.8721 0.244 0.8397 0.9888 0.9976
Table 1: Results of quality of text generate in PEGS and baseline.
5.1 Experimental Settings
Implementation Details
At each training stage, the AdamW optimizer is adopted, and the cosine annealing scheduler is used to adjust the learning rate with the initial learning rate of 1e-4. The details of pre-training have been introduced in Section 4. Our model is fine-tuned on the entire StickerConv. It is noteworthy that only 3% of the model parameters are updated, and it takes about 6 hours on a single A800 GPU.
Baseline
We compare with the following baselines: (1) Vicuna-7B: An open-source chatbot trained by fine-tuning LLaMA on user-shared conversations collected from ShareGPT Chiang et al. (2023). We use the Vicuna-delta-v0 version to ensure consistency with the model base used during training. (2) ChatGLM3-6B: An open-source, bilingual supported large-scale dialogue model Zeng et al. (2023).
5.2 Evaluation Metrics
5.2.1 General Metrics
This category focuses on assessing the quality of the model’s textual output.
Diversity
We adopt Distinct-n (Dist-1, 2, 3) Li et al. (2016) as quantitative indicators of diversity, measuring the variety and creativity of the model’s outputs.
Empathy Performance
To evaluate the model’s capacity for empathy, we adopt the EPITOME Sharma et al. (2020a) method, comprising three sub-metrics: IP (interpretations), EX (explorations) and ER (emotional reactions).
Empathy Difference
DIFF-EPITOME Lee et al. (2022) measures the difference scores of IP, EX, ER between the golden response and predicted response.
5.2.2 LLM-based Metrics
LLM has been widely proven to be capable of acting as a grader similar to humans and can provide reasonable scores based on the designed prompts. These metrics not only focues on textual outputs but also include sticker outputs, forming a comprehensive multimodal evaluation system.
Empathy
Utilizing ChatGPT-3.5-turbo as a scorer, based on the context, we assess empathy scores for the model’s textual and sticker responses, denoted as Emp-txt and Emp-sticker, respectively. To mitigate randomness and bias, we conduct five independent scorings for each response and calculate the average.
Consistency
Again using ChatGPT-3.5-turbo as a scorer, based on the context, we assign consistency scores for textual and sticker responses, marked as Con-txt and Con-sticker, employing the same scoring method.
Ranking
Employing ChatGPT-3.5-turbo as a result ranker, we compare responses of different models to the same context. The ranker organizes responses based on quality, empathy, and consistency, averaging across all possible ranking combinations to ensure fairness and objectivity in assessment.
5.2.3 Multimodal Metrics
CLIP-Score
We utilize CLIP (Radford et al., 2021) as multimodal encoder, evaluate the consistency of model’s multimodal responses.
MM-Relevance
MM-Relevance Feng et al. (2023) is a metric designed to measure the relevance of multimodal responses. Traditional mm-relevance metrics, however, face limitations, particularly in cases where there is a lack of alignment between model responses and golden responses. To address these challenges, we propose Inter-MM-Rel (Interactive multimodal Relevance). This innovative approach calculates and averages the similarity scores between each mode of information, without the constraint of alignment, allowing for a more comprehensive and accurate assessment of multimodal response relevance.
5.3 Result Analysis
From Table 1, PEGS shows remarkable performance in several metrics, particularly in EX (Explorations), DIFF-EPITOME and diversity measuresm indicating a superior ability in generating diverse and exploratory responses. However, PEGS’s lower scores in IP, ER could be due to two reasons: (1) The general metrics only evaluate the text output and underestimate the actual empathy ability of the PEGS. (2) In multimodal empathy dialogue task, stickers play a significant role in conveying emotional reactions and explanations, while text handles more explorative tasks.
Models Rank Emp-txt Emp-sticker Con-txt Con-sticker
Vicuna 2.12 3.6768 4.3221
ChatGLM3 1.99 3.6908 4.3405
PEGS 1.88 3.7354 4.2534 4.2739 4.1155
Table 2: Results of LLM-Scorer.
Table 2 reveals PEGS’s strong performance in ranking, empathy text (Emp-txt), empathy sticker, and consistency of sticker responses (Con-sticker). However, its lower consistency in text (Con-txt) might be attributed to its emphasis on empathetic stickers, which could compromise some textual context consistency. This is in line with the additional information suggesting that the empathetic effectiveness of stickers might require some trade-offs in textual context consistency.
Models MM-Rel Inter-MM-Rel CLIP-Score
PEGS 69.41 59.06 21.03
Table 3: Results of multimodal evaluate.
Table 3 showcases the multimodal capabilities of PEGS. PEGS performs well in multimodal relevance (MM-Rel and Inter-MM-Rel) and consistency (CLIP-Score). This suggests that PEGS effectively integrates different modalities in its responses, maintaining relevance and consistency across them.
6 Conclusion
We explored the concept of multimodal empathetic response and created the StickerConv, the first dataset specifically designed for multimodal empathetic dialogue. We developed the Agent for StickerConv, a sophisticated LLM-based multi-agent system capable of simulating human-like interactions using stickers, thereby creating multimodal empathetic responses. Building upon the StickerConv, we developed PErceive and Generate Stickers (PEGS), an advanced multimodal dialogue model. This model is adept at both perceiving and generating stickers, effectively enhancing the communication experience. Furthermore, we established a comprehensive set of evaluation metrics for multimodal empathy tasks based on LLM. We are confident this work will be a valuable asset in advancing research in the field of multimodal empathetic dialogue systems.
Appendix A Appendix
A.1 Dataset Statistics
Table 4 offers a statistical breakdown of the dataset, segmented into training, validation, and test splits. It provides a quantitative overview, detailing the number of instances in each subset, the unique stickers, turn counts, and token counts.
split Number Unique Sticker Turn $\text{Token}_{\text{vicuna}}$ $\text{Token}_{\text{GPT}}$
train 10,785 4,798 59,424 10,681,108 9,070,221
validation 1,000 880 5,496 997,569 845,759
test 1,146 1,439 6,128 659,695 773,473
Table 4: The statistics of StickerConv.
Figure 5 showcases the emotion distribution within user profiles in the Agent for StickerConv. The chart illustrates the percentage share of each emotion, offering insight into the prevalence of affective expressions within the dataset.
Figure 5: Emotion distribution of user profile in Agent for StickerConv.
Figure 6, a word cloud, visualizes the 200 most prevalent emotion-related words extracted from the StickerConv. The prominence of each term in the cloud is indicative of its frequency within the dataset. This figure serves to underscore the linguistic diversity and the emotional range encapsulated in the dataset.
Figure 6: The 200 most popular emotion-related words in StickerConv.
Appendix B Case Study
We conduct a series of case studies to provide a more straightforward illustration of the performance of our model instruction-tuned by the StickerConv in multimodal empatheric conversations. These cases demonstrate the ability of PEGS to empathize with users and perceive the stickers provided by users, along with generating multimodal responses with vivid stickers.
As shown in Figure 7, in the positive emotion condition, PEGS will reward users with praise and support. As shown in Figure 8, two conversations in a negative scenario are displayed. PEGS as a conciliator can comfort users and guide their emotions towards positivity.
Appendix C Prompts
All {format_instructions} in the prompt utilize the response of langchain [^2] https://www.langchain.com/’s response_schema function implements format control. For detailed formatting information, please refer to our project’s open-source repository: https://github.com/ZhangYiqun018/StickerConv.
C.1 Sticker Process
Figure 9 shows the prompt of sticker process. For each sticker, we subject it to a process of filtering, annotation, and knowledge extraction, ultimately yielding three pieces of information: description, emotion and recommendation.
C.2 Sticker Agent Chat
Figure 10 shows the chat prompt, the summary encompasses the persona of the user or system, the current status, and the core traits generated based on memory. Figure 11, 12, 13 show the process of used sticker.
C.3 LLM-Based Scorer
Figure 14 illustrates the template for the preference rank scorer. This scorer integrates the outputs from all baseline models as well as PEGS, constituting the {response_list}. It then proceeds to rank these responses and provide reasoning for the assigned rankings.
Figures 15 and 16 depict the templates for the consistency scorer. The selection of the appropriate template depends on the modality of the response generated by the model. Specifically, Figure 15 is utilized for scoring purely textual responses. In contrast, when the response is multimodal, the scoring is conducted using the template shown in Figure 16.
Similarly, Figures 17 and 18 present the templates for the empathy scorer. The methodology for processing responses of different modalities aligns with the approach described for the consistency scorer.
C.4 Joint Learning
Figure 19 shows the instruction prompts for joint learning. We employed an instruction template similar to Vicuna_delta_v0. For every query or response that uses a sticker, we append an <IMG> identifier at the end.
Figure 7: Examples of conversations with positive emotions. Users can chat with multimodal content (text and stickers) and will receive positive multimodal responses.
Figure 8: Examples of conversations with negative emotions. Our model can empathize with users who are suffering from negative emotions, such as sadness (left) or anger (right) . Users will be comforted and guided to positive sentiments.
Figure 9: Prompt Templates for Process Sticker.
Figure 10: Prompt Template for Chat Prompt.
Figure 11: Prompt Template for Sticker Intent.
Figure 12: Prompt Template for Sticker Query.
Figure 13: Prompt Template for Sticker Select.
Figure 14: Prompt Template for Ranking score.
Figure 15: Prompt Template for Consistency Scorer (text only).
Figure 16: Prompt Template for Consistency Scorer (multimodal).
Figure 17: Prompt Template for Empathy Scorer (text only).
Figure 18: Prompt Template for Empathy Scorer (multimodal).
Figure 19: Prompt Template for Joint Learning.

### 16 Feb 2024 v2

@@
 Abstract
-Stickers, while widely recognized for enhancing empathetic communication in online interactions, remain underexplored in current empathetic dialogue research.
+Stickers, while widely recognized for enhancing empathetic communication in online interactions, remain underexplored in current empathetic dialogue research, notably due to the challenge of a lack of comprehensive datasets.
 In this paper, we introduce the Agent for STICKERCONV (Agent4SC), which uses collaborative agent interactions to realistically simulate human behavior with sticker usage, thereby enhancing multimodal empathetic communication.
-Building on this foundation, we develop a multimodal empathetic dialogue dataset, StickerConv, which includes 12.9K dialogue sessions, 5.8K unique stickers, and 2K diverse conversational scenarios, specifically designs to augment the generation of empathetic responses in a multimodal context.
-To leverage the richness of this dataset, we propose PErceive and Generate Stickers (PEGS), a multimodal empathetic response generation model, complemented by a comprehensive set of empathy evaluation metrics based on LLM.
-Our experiments demonstrate PEGS’s effectiveness in generating contextually relevant and emotionally resonant multimodal empathetic responses, contributing to the advancement of more nuanced and engaging empathetic dialogue systems.
-Our project page is available at https://neu-datamining.github.io/StickerConv.
-StickerConv: Generating Multimodal Empathetic Responses from Scratch
+Building on this foundation, we develop a multimodal empathetic dialogue dataset, STICKERCONV, comprising 12.9K dialogue sessions, 5.8K unique stickers, and 2K diverse conversational scenarios.
+This dataset serves as a benchmark for multimodal empathetic generation.
+To advance further, we propose PErceive and Generate Stickers (PEGS), a multimodal empathetic response generation framework, complemented by a comprehensive set of empathy evaluation metrics based on LLM.
+Our experiments demonstrate PEGS’s effectiveness in generating contextually relevant and emotionally resonant multimodal empathetic responses, contributing to the advancement of more nuanced and engaging empathetic dialogue systems [^1] https://github.com/ZhangYiqun018/StickerConv.
+STICKERCONV: Generating Multimodal Empathetic Responses
+from Scratch
 […]
@@
 Both parties can utilize the stickers to express their emotions, which enhances interactivity and expression.
-Assistant can empathize with the user according to the conversation (green text).
-Acknowledging stickers as a crucial non-textual modality and incorporating them into empathetic response generation offers novel perspectives and methodologies for empathetic dialogue tasks Zech and Rimé (2005); Rashkin et al.
-(2019).
-Specifically, integrating text and stickers paves the way for a new multimodal empathetic communication approach, targeting the generation of more diverse and high-quality empathetic responses.
+The assistant can empathize with the user according to the conversation (green text).
+Integrating stickers with textual communication and interspersing stickers within the dialogue can yield more varied and superior-quality empathetic replies.
 A primary challenge in integrating stickers into empathetic response generation is developing a high-quality dataset to support this innovative multimodal communication.
-Traditional dataset construction methods, including expensive human crowdsourcing, privacy issues Xu et al.
-(2022), the difficulty in acquiring data from social software companies, and the variable quality of Internet-crawled data Foerderer (2023), face significant limitations.
-To address this, leveraging large language models (LLMs) for dataset construction is proposed.
+To address this, we leverage large language model (LLMs) for dataset construction.
 LLMs, with their extensive world knowledge and text processing capabilities (OpenAI, 2023; Touvron et al., 2023a, b), demonstrate near-human annotation abilities Wang et al.
@@
 (2023).
-However, LLMs have limitations in empathetic tasks, excelling in responding to explicit human instructions but lacking in proactivity Wu et al.
-(2023a); Yin et al.
-(2023), a critical aspect of empathy.
+However, applying LLMs directly have limitations in empathetic tasks, excelling in responding to explicit human instructions but lacking proactivity, a critical aspect of empathy Wu et al.
+(2023); Yin et al.
+(2023).
 Empathy necessitates understanding others’ emotions and the ability to actively express support and understanding Zech and Rimé (2005); Sharma et al.
-(2020a).
+(2020).
 To mitigate this, we introduce a multi-agent system based on LLMs, Agent for STICKERCONV (Agent4SC).
@@
 It not only generates text responses but also strategically selects suitable stickers, thereby effectively enhancing empathy.
-Based on the superior capabilities and interactions of Agent4SC, we built a multimodal empathetic dataset, StickerConv, that comprise 12.9K dialogue sessions and 5.8K unique stickers.
-StickerConv boasts an average of 5.22 stickers per dialogue session.
+Based on Agent4SC, we build a multimodal empathetic dataset, STICKERCONV, that comprises 12.9K dialogue sessions and 5.8K unique stickers.
+STICKERCONV boasts an average of 5.22 stickers per dialogue session, mirroring the sticker usage patterns observed in human communication.
 Figure 1 depicts an example of conversations in our dataset.
-To the best of out knowledge, this is the first multimodal empathetic dialogue dataset, with the particular utility of sticker as non-textual modal information to better facilitate empathy.
-Although Agent4SC exhibits sound multimodal empathetic dialog response generation capability, it is limited by expensive inference costs, while the retrieval approach leads to poor performance in new scenarios.
-To further advance the research on multimodal empathetic dialog, based on StickerConv, we develop a end-to-end large multimodal dialog model, PEGS, with the ability to PErceive and Generate Stickers, and has a significantly lower deployment and inference cost.
+To the best of our knowledge, this is the first multimodal empathetic dialogue dataset, with the particular utility of sticker as non-textual modal information to better facilitate empathy.
+Although Agent4SC effectively generates multimodal empathetic responses, it is limited by expensive inference costs and specific sticker databases.
+To further advance the research on multimodal empathetic dialogue, we develop an end-to-end multimodal empathetic response generation framework, PEGS, with the ability to PErceive and Generate Stickers.
 Beyond the general ability to generate textual empathetic responses, PEGS receives multimodal inputs and autonomously generates stickers based on the emotional and contextual aspects of the dialogue at the appropriate moment.
 Furthermore, to simulate human communications on social media in the real world, our model supports interleaved multiple image and text inputs.
-Due to potential misalignments between the modality quantities of predicted responses and those of the golden responses, reliance on a singular modal metric may lead to significant distortions in evaluation results.
-Empathy, being inherently subjective, presents challenges in its direct quantification.
-In this context, we introduce a novel multimodal empathetic response evaluation method.
-This method enables the integration of multimodal information, emphasizing three critical dimensions: empathy, consistency, and preference ranking.
-Leveraging the world knowledge and anthropomorphic capabilities of LLMs, our method offers robust support for evaluating multimodal empathetic responses.
+Misalignments between the modal quantities of predicted and golden responses can distort evaluation outcomes, and empathy is difficult to quantify due to its subjective nature.
+To address this, we propose a novel method for evaluating multimodal empathetic responses, focusing on empathy, consistency, and ranking.
+Utilizing the extensive world knowledge and anthropomorphic abilities of LLMs, this approach provides solid support for assessing multimodal empathetic replies.
 In conclusion, the main contributions of this work are as follows:
-  • We introduce Agent for StickerConv (Agent4SC), a LLM-based multi-agent system, integrates stickers intelligently into empatheric dialogues.
-Agent4SC ensures the stickers are contextually consistent, varied, and empathetically aligned with human interaction.
-Leveraging the Agent4SC, we built a multimodal empatheric dialogue dataset, StickerConv, comprising 12.9K dialogue sessions and 5.8K unique stickers.
-  • We design PErceive and Generate Stickers (PEGS), an empathetic dialogue response generation model that, within its end-to-end framework, intuitively incorporates stickers based on the emotional and contextual dynamics of a dialogue.
-PEGS adeptly processes multimodal inputs, generating empathetic textual responses and appropriately using stickers to enrich these replies.
-  • We propose a multimodal empathetic response evaluation approach.
-This method integrates multimodal information and concentrates on three principal dimensions: empathy, consistency, and preference ranking.
+  • We introduce an LLM-based multi-agent system, Agent for STICKERCONV (Agent4SC), which integrates stickers into empathetic dialogues, ensuring contextual consistency, variety, and empathy aligned with human interactions.
+Using Agent4SC, we create a multimodal empathetic dialogue dataset, STICKERCONV.
+  • We design PErceive and Generate Stickers (PEGS), a multimodal empathetic response generation framework that intuitively incorporates stickers based on the emotional and contextual dynamics of the dialogue.
+PEGS adeptly processes multimodal inputs, generating empathetic textual responses and using stickers appropriately to enhance these responses.
+  • We propose a method for assessing multimodal empathetic responses. It leverages LLM to evaluate the quality of these responses, with a specific focus on empathy, consistency, and ranking.
 2 Related Work
 2.1 Empathetic Response Generation
-The purpose of empathetic response generation is to enable machines to understand and respond to human emotions and feelings.
-In this field, The creation of the EMPATHETICDIALOGUES dataset Rashkin et al.
-(2019) laid the foundation.
-Subsequently, the empathetic listener mixture model Lin et al.
-(2019), which emphasizes understanding and appropriately responding to different emotions, and studies focusing on emotion imitation and emotion cause words to enhance the effectiveness of empathetic dialogue Majumder et al.
-(2020) have achieved notable results.
-Additionally, the introduction of the EPITOME conceptual framework Sharma et al.
-(2020b), a classification system for empathetic response intents Welivita and Pu (2020), research on generating empathetic responses based on commensense knowledge Sabour et al.
-(2021), and the proposal of the multifactorial hierarchical framework CoMAE Zheng et al.
-(2021) have provided new perspectives and methods for the field.
-Despite these advancements, the research in this domain is not without its limitations.
-Primarily, the field has been domainated by small-scale models, with limited application of large language models, such as ChatGPT OpenAI (2023), in the generation of empathetic responses Lee et al.
+Empathetic response generation focuses on enabling machines to understand and respond to human emotions.
+The foundational EMPATHETICDIALOGUES dataset Rashkin et al.
+(2019) and subsequent innovations like the empathetic listener mixture model Lin et al.
+(2019) have significantly advanced this area.
+Large Language Models (LLMs) like ChatGPT have been explored for empathetic response generation, although their application remains limited Lee et al.
 (2022); Zhou et al.
 (2023b).
-Secondly, the focus has predominantly been on text-based empathetic responses, thus neglecting the rich potential of multimodal information, such as image.
-Emerging studies underscore the significance of such multimodal data, especially stickers, in alleviating stress, enhancing well-being, and fostering emotional resonance and social connectivity Kariko and Anasih (2019); Akram and Drabble (2022); Jiang et al.
-(2023).
-Finally, the subjectivity and multidimensionality of empathetic tasks make it challenging to quantitatively assess their effectiveness Fu et al.
+Challenges persist, especially in leveraging multimodal information for richer emotional engagement and in accurately evaluating empathetic responses due to their subjective nature.
+These obstacles highlight the ongoing need for research in effectively integrating LLMs and multimodal data into empathetic dialogue.
+Additionally, the subjective nature of empathy complicates its quantitative assessment Fu et al.
 (2023); Lee et al.
-(2022), especially in the case of multimodal empathetic responses, posing a significant challenge in the field of empathetic response generation.
+(2022), posing a further obstacle to the field’s advancement.
 2.2 Large Multimodal Models
-Multimodal Perception
-Large language models (LLMs), such as ChatGPT (OpenAI, 2023), LLaMA (Touvron et al., 2023a, b), demonstrate powerful capabilities in dialog interaction and instruction following, and recent researches have extended LLMs to multimodal domains.
+LLMs, such as ChatGPT (OpenAI, 2023), LLaMA (Touvron et al., 2023a, b), demonstrate powerful capabilities in dialog interaction and instruction following, and recent researches have extended LLMs to multimodal domains.
 Flamingo (Alayrac et al., 2022) exhibits promising zero-shot and few-shot multimodal understanding capability by adding a cross-attention layer to connect the frozen vision encoder with the LLM.
 BLIP (Li et al., 2023b; Dai et al., 2023), MiniGPT-4 (Zhu et al., 2023; Chen et al., 2023) and LLaVA (Liu et al., 2023b, a) bridge the frozen vision encoder and the LLM through a small intermediate model.
-Multimodal Generation
-The birth and development of the Diffusion Models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Rombach et al., 2022) significantly contribute to the evolution of image generation.
 Koh et al.
 (2023) for the first time explore the mapping of the output of LLMs into the input space of the vision decoder, empowering LLMs with image generation capability.
-GILL (Koh et al., 2023), NExT-GPT (Wu et al., 2023b), MiniGPT-5 (Zheng et al., 2023) align LLMs to frozen vision decoders through an encoder-decoder transformer, while SEED (Ge et al., 2023) employs a learnable Q-Former.
+GILL (Koh et al., 2023), MiniGPT-5 (Zheng et al., 2023) align LLMs to frozen vision decoders through an encoder-decoder transformer, while SEED (Ge et al., 2023) employs a learnable Q-Former.
 In contrast to previous works which favor realistic images, our target is to generate stickers, which are abstract and exhibit distinct emotional tendencies.
 2.3 LLM-Based Agents
-In the realm of artificial intelligence, Large Language Model (LLM)-based agents epitomize a cutting-edge paradigm, leveraging the robust computational prowess of LLMs.
-These agents are not merely confined to problem-solving and task execution; their capabilities extend to intricate reasoning and sophisticated interactions, as elucidated in recent studies Wang et al.
+LLM-based agents mark a major leap in AI, leveraging their capabilities for tasks like reasoning and interaction, as shown in recent studies Wang et al.
 (2023b); Li et al.
 (2023a).
-The versatility of LLM agents is further demonstrated through their adaptability in diverse domains such as software development Qian et al.
-(2023) and scientific research Boiko et al.
-(2023).
-A remarkable aspect of LLM-based agents is their ability to emulate complex human social behaviors.
-This includes a spectrum of interactions ranging from social engagements Park et al.
+They find uses across various domains, such as software engineering Qian et al.
+(2023) and scientific inquiry Boiko et al.
+(2023), highlighting their versatility.
+These agents can imitate complex human actions, partake in social interactions Park et al.
 (2023); Tu et al.
-(2023), to simulating multifaceted scenarios like elections Argyle et al.
+(2023), and replicate intricate scenarios like elections Argyle et al.
 (2022), debates Wang et al.
-(2023a), and consumer activities Wang et al.
-(2023c), Such capabilities underscore the potential of LLM agents in understanding and replicating human social dynamics.
-The most transformative application of LLM-based agents, however, lies in their utility for generating training data for diverse AI models.
+(2023a), and consumer patterns Wang et al.
+(2023c), illustrating their capacity to emulate human social dynamics.
+A notable innovation is their use in generating AI training data.
 Studies Wang et al.
@@
 (2023); Zhou et al.
-(2023a) have been instrumental in showcasing the efficacy of these agents in producing high-quality data, paralleling human annotators, yet at a substantially lower cost.
-This breakthrough has redefined the paradigm for creating training datasets, offering a more practical and cost-effective solution in the field of AI model development.
-Figure 2: The Overview of Agent for StickerConv.
+(2023a) highlight their efficiency and cost-effectiveness in producing high-quality training datasets, transforming AI model development.
+Figure 2: The overview of Agent4SC.
+Memory and Plan modules enable the agent to mimic human observation and thought, overcoming LLMs’ inability to grasp nuanced emotions.
+The Action module supports generating insights with human-like emotional reactions.
+The Profile module gives each agent distinct reflections and actions.
+Furthermore, Agent4SC uses stickers as a Tool for more natural conversation, allowing the agent to choose stickers like humans.
+These modules streamline observation, reflection, and action, while the Manager Agent maintains performance and quality.
 3 Agent for STICKERCONV
-The limitations of Large Language Models (LLMs) and Language Model Multimodalities (LMMs), are twofold: (1) They struggle with capturing nuanced human emotions and providing empathetic responses.
-(2) While adept at responding to explicit human instructions, they lack the initiative to act beyond these directives.
-This inherent disadvantage hinders their performance in tasks that require empathy, making them less effective in this area compared to others.
-To bridge this gap, we propose the Agent for StickerConv (Agent4SC), a novel LLM-based multi-agent.
-This system is designed to simulate human dialogue by leveraging the collaborative efforts of multiple modules, along with the use of stickers, to produce emotional and diverse empathetic responses.
-The architecture of Agent4SC is depicted in Figure 2 and comprises two key agents.
-The User Agent, forming the core of the system, is an intricate assembly of several modules.
-It replicates the human processes of observation and reflection through the interplay between its Memory and Plan modules.
-This design effectively addresses the LLMs’ shortcomings in comprehending subtle human emotions.
-The insights gained from observation and reflection are articulated via the Action module.
-Consequently, the Agent is more adept at generating human-like emotional responses.
-The entire process is orchestrated by the Profile module, which imparts distinctiveness in reflection and behavior to different User Agents.
-Furthermore, Agent4SC adopts a pioneering approach by treating stickers as a Tool of communication.
-Agents autonomously select appropriate stickers for use, mirroring human behavior in conversations.
-This function is facilitated through the integrated process of observation, reflection, and action.
-The User Agent ensures consistent performance and quality control with the support of the Manager Agent.
-Agent4SC is not just a static tool but a dynamic participant in conversations, capable of grasping the subtleties of human emotions and the appropriate moments for empathetic interaction.
+Confronted with the pivotal challenge of a lack of datasets for multimodal empathetic response tasks, we craft our own dataset utilizing large language models (LLMs).
+Nevertheless, LLMs experience difficulties in grasping nuanced human emotions and initiating actions beyond explicit directives.
+These limitations render LLMs and large multimodal models (LMMs) less proficient in the generation of multimodal empathetic responses.
+In response to these issues, we introduce Agent for STICKERCONV (Agent4SC), a multi-agent system predicated on LLM, devised to mimic human conversational patterns.
+Figure 2 presents the overview of Agent4SC.
+By integrating multiple modules and the strategic use of stickers, Agent4SC aims to generate emotional and varied empathetic responses, thereby overcoming the inherent deficiencies of LLMs in empathetic engagements.
 3.1 Profile Module
-The profile module serves as the foundation of the Agent4SC, shaping the personality and behavioral patterns of users in empathetic interactions.
-It comprises two components: Persona and Situation.
-The Persona section vividly portrays the user’s character traits, life background, and experiences, whereas the Situation part focuses on the user’s current circumstances, emotional state, and potential challenges they may encounter.
-To develop rich and diversified user profiles, we initially manually crafted profiles with a variety of emotioinal distributions, subsequently expanding them to 2K unique profiles through the SELF-INSTRUCT Wang et al.
-(2023d) method by ChatGPT-3.5-turbo.
-Our constructed profile archive features a comprehensive emotional distribution, as illustrated in Figure 5.
-We believe that emotionally rich and content-diverse user profiles will enable the agent system to more authentically simulate human responses in empathetic dialogues.
+The profile module, comprising Persona and Situation components, underpins its foundation by defining users’ personality traits and behavioral patterns for empathetic interactions.
+Persona outlines users’ character traits, backgrounds, and experiences, while Situation details their current circumstances and emotional states.
+To enrich and diversify user profiles, we initially create profiles with varied emotional distributions, then expand to 2,000 unique profiles using SELF-INSTRUCT Wang et al.
+(2023d) method.
+This approach results in a profile archive with a broad emotional spectrum, as shown in Figure 5, aiming to enhance the system’s ability to simulate human-like responses in empathetic dialogues.
 3.2 Tool Module
-To effectively simulate human behavior in using stickers for communication, we developed a Tool Module, converting the SER30K dataset into a functional tool.
-This dataset comprises 1,887 themes and 30,739 emotion-labeled stickers (Liu et al., 2022), such as happiness, anger, neutral, disgust, fear, sadness and surprise.
-In integrating it with the agent system, we encountered several challenges: (1) The dataset contains a significant number of stickers unsuitable for casual conversation or empathetic communication.
-(2) Stickers are annotated with a single emotional tag from a set of seven, lacking comprehensive content descriptions and a detailed emotional analysis.
-This limitation hinders the agent’s ability to appropriately utilize stickers based solely on these tags.
-(3) Humans selecting stickers in conversation typically do not choose from a database containing tens of thousands of sticker.
-Instead, they prefer to pick from a customized and smaller-scale sticker library.
-To overcome these challenges, we design a series of sophisticated sticker processes to refashion sticker into an effective tool for accurate agent application:
-Filter
-Leveraging the LLaVa-v1.5-13b (Liu et al., 2023a), we crafted a nuanced three-step Chain of Thought Wei et al.
-(2023) procedure to meticulously scrutinize each sticker in SER30K, as shown in Figure 9.
-This approach entailed the model initially summarizing the sticker’s content, subsequently assessing its relevance to casual conversation, and finally determining suitability through iterative responses.
-This strategy not only affirmed selection accuracy but also mitigated model-induced randomness and hallucination discrepancies.
-Annotation
-For filtered stickers, a second iteration of the CoT process was executed using the same model.
-This phase involved responding to five pivotal queries, covering aspects like the sticker’s content, emotional conveyance, applicable scenarios, emotional tone, and potential satirical or humorous connotations (shown in Figure 9).
-These responses yielded comprehensive insights into each sticker, subsequently formatted into Q&A pairs, awaiting further processing.
-Knowledge
-Post-acquisition of intricate sticker details, the ChatGPT-3.5-Turbo model was engaged for an additional processing phase.
-Relying on the established Q&A pairs, this model adeptly deduced each sticker’s specific contextual suitability, culminating in the distillation of three fundamental knowledge: description, emotion, recommendation.
-These synthesized insights are instrumental in enriching agent’s understanding of sticker, and enable them to apply stickers correctly.
-Split
-In this phase of refining the SER30K sticker dataset for integration into our agent system, we focus on dividing the database into smaller, more manageable subsets.
-Each subset is transformed into a vector database, where the embeddings represent a unique blend of the original emotional labels of the stickers and the enriched emotional knowledge extracted in the previous steps.
-After the above-mentioned processes, the original SER30K dataset has been transformed into a series of refined, smaller-scale vector databases.
-However, a crucial step is still required to make it a directly usable tool for agents.
-To address this, we have defined a specific module operation, termed "Retrieve".
-In this process, when an agent intends to use a sticker, it queries its accessible sub-vector database with its intended use and desired emotional expression of the sticker as search criteria.
-The system then retrieves and returns the top-K stickers with the highest similarity.
+To adapt the SER30K dataset, with its 1,887 themes and 30,739 emotion-tagged stickers (Liu et al., 2022), for human-like sticker use in dialogues, we transform it into a tool.
+Integrating it into the agent system faced several hurdles: the inappropriateness of many stickers for empathetic dialogues, the lack of detailed content or emotional analysis in the single-emotion labeled stickers, and the contrast between the extensive SER30K collection and the limited, personalized collections humans use.
+To make stickers an effective tool, we refine the SER30K through filtering, annotating, and knowledge extraction (detailed in Appendix B.1).
+Each sticker is enriched with three pieces of information: emotion, description, and recommendation.
+Subsequently, SER30K is segmented into smaller, emotionally balanced vector databases.
+A critical action, "Retrieve", is then designed for agents to efficiently use these databases by searching for stickers based on the communication context and desired emotion, with the system suggesting the top-K closest matches for selection.
+3.3 Memory Module
+In empathetic interactions, response quality hinges on communication coherence and personality trait consistency.
+Agents must learn and give feedback by observing during interactions and ensure long-term personality consistency.
+Following the Generative Agent Park et al.
+(2023), we use short-term updates of core traits from profiles and observations for immediate memory.
+For long-term memory, continuous interactions and self-reflection help maintain consistency.
 3.4 Plan Module
-The Plan module is a core component of the Agent4SC, responsible for orchestrating the behavior of agents in empathetic interactions.
-To ensure that agents mimic human behavior in using stickers, we have specially designed three key actions to ensure the timely use of stickers, their coherence with the context, and their effective role in empathy.
-  • Intention: Under the control of Profile module, the agent assesses the potential for using emotive stickers based on interaction history and observational data.
-The prompt template for this action is shown in Figure 11.
-  • Query: If the agent intends to use a sticker, it will describe the emotional content of the sticker in detail.
-This description will be influenced by the Profile module, observation, and the specific intent behind using the sticker.
-The prompt template for this action is shown in Figure 12.
-  • Select: The agent sends the query generated in the previous step to the sticker database accessible to it, retrieving the top-K stickers with the highest similarity.
-Subsequently, the agent combines the candidated sticker’s index, description, emotion and recommendation into a list.
-From this list, the agent selects the sticker it wishes to use.
-The prompt template for this action is shown in Figure 13.
-3.4 Memory Module
-In empathetic interactions, the coherence of communication and the consistency of personality traits are fundamental to the quality of responses.
-This necessitates that agents not only learn and provide feedback through observation during the interaction process, but also maintain long-term consistency in their personality traits.
-We adopted a strategy similar to that of Generative Agent Park et al.
-(2023): for short-term memory, we update the agent’s core traits based on the profile and observations; for long-term memory, the agent combines continuous interaction with external entities and regular self-reflection to maintain consistency.
+The Plan module orchestrates agent behavior in empathetic interactions.
+To ensure that agents mimic human behavior in using stickers, we specially design three key actions to ensure the timely use of stickers, their coherence with the context, and their effective role in empathy.
+  • Intention: Directed by the Profile module, the agent assesses the use of emotive stickers, considering interaction history and observed data.
+  • Query: When intending to use a sticker, the agent describes its emotion, shaped by the Profile module, observations, and the intent for its use.
+  • Select: After querying its sticker database with the crafted query, the agent retrieves the top-K most relevant stickers.
+It organizes these stickers into a list detailing their emotion, description, and usage recommendation, from which it selects the most appropriate sticker.
 3.5 Action Module
@@
 It is responsible for receiving instructions and inputs from other parts of the system and executes two main actions based on this information:
-  • Message: generates text responses based on the Agent’s Profile and Memory module.
-  • Sticker: engages in sticker usage as dictated by the Plan module.
+  • Message: The agent generates text responses based on the agent’s Profile and Memory module.
+  • Sticker: The agent engages in sticker usage as dictated by the Plan module.
 3.6 Manager Agent
-The Manager Agent, a critical componet in enhancing the performance of the Agent4SC, plays a significant role in ensuring interaction quality and dialogue consistency.
-This agent comprises two main modules:
-Quality Assurance
-This module is crucial for upholding high dialogue standards, It accomplishes this by: (1) Evaluating the relevance of stickers within the conversation, ensuring that they contribute meaningfully to the dialogue.
-(2) Regulating the use of stickers to prevent their overuse, thus maintaining the integrity and flow of the conversation.
-(3) Performing regular checks on dialogue quality, with a special focus on the appropriateness of response length and the avoidance of redundant information.
-Process Control
-This module takes charge of guiding the dialogue’s progression. It leverages historical interactions as a reference to make informed decisions on the optimal timing for concluding dialogues.
-Figure 3: The chart of emotional distribution in the choice of stickers between users and the system.
-This chart revealed a striking trend: users have a significant preference for stickers that convey negative emotions, in contrast to the system.
-The system predominantly utilizes stickers expressing neutral and positive emotions.
-This comparison not only reflects the distinct emotional expression preferences between users and the system but also highlights the system’s active and supportive role in interactions.
+The Manager Agent, a critical component in enhancing the performance of the Agent4SC, plays a significant role in ensuring interaction quality and dialogue consistency.
+This agent comprises two main modules: (1) Quality Assurance: This module plays a key role in maintaining high dialogue standards by evaluating sticker relevance to enrich conversations, regulating sticker usage to avoid misuse, and conducting quality checks focused on response length appropriateness and redundancy avoidance.
+(2) Process Control: This module takes charge of guiding the dialogue’s progression.
+It leverages historical interactions as a reference to make informed decisions on the optimal timing for concluding dialogues.
 3.7 The STICKERCONV Dataset
-The Agent4SC facilitates empathetic interaction with humans through the collaborative interplay of its modules, and possesses the capability to utilize stickers.
-Leveraging this, we propose the construction of a large-scale multimodal empathetic dialogue dataset, StickerConv, with stickers serving as a non-textual modality.
-Figure 4: The overview architecture of PEGS.
-The input images are jointly encoded by image encoder, Q-Former and a linear layer.
-Vicuna is used as the language model.
-To empower the language model to generate images, the output of the LLM is first mapped into the input space of the image decoder through a feature mapper, and then the frozen image decoder is employed to generate images.
-In the dataset creation process, two roles are designated: User and System.
-The User, acting as the dialogue initiator, is assigned a profile randomly selected from 2,000 personality profiles generated as described in Section 3.1.
-The System, in the roles of a listener and empathizer, is assigned a predefined positive profile.
-Regarding the sticker tool, we have processed and subdivided the sticker dataset into 100 smaller vector databases.
-In each dialogue session, both User and System have access to one randomly assigned vector database from this pool.
-This arrangement is designed to better simulate human sticker usage behavior while enhancing the diversity and richness of stickers.
-It is important to note that vector databases numbered 1 to 80 are exclusively employed for the creation of training and cross-validation sets, whereas those numbered 81 to 100 are reserved for the construction of the test set.
-StickerConv consists of 12,931 dialogue sessions, 67,505 stickers (unique 5.8K stickers) and over 2K conversation topic.
-The detailed statistical information of our dataset is shown in Table 4.
-On average, each dialogue session in StickerConv has 5.22 stickers and 5.49 turns.
-The emotional label distribution of the stickers used by User and System is shown in Figure 3.
-This analysis reveals notable differences in the emotional distribution of sticker usage, attributable to the distinct roles of the User and the System.
+Leveraging Agent4SC, we propose the construction of a multimodal empathetic dialogue dataset, STICKERCONV.
+Figure 3: The architecture of PEGS framework includes various routing options, distinguished by colored connecting lines.
+Input stickers undergo joint encoding by an image encoder, Q-Former, and a linear layer, with Vicuna serving as the language model.
+The output of the LLM activates two sets of tokens differently across model versions: one for image retrieval and the other as a textual condition.
+Subsequently, the frozen image decoder generates images.
+In our dataset creation, we define two roles: User, the dialogue initiator with a profile from 2,000 generated personalities, and System, acting as a listener and empathizer.
+To mimic the human use of stickers, the sticker dataset is divided into 100 vector databases, each ensuring a consistent emotional distribution of stickers.
+The top-K of Tool module is set to 10.
+For each dialogue session, both User and System access a randomly chosen vector database, mirroring human sticker usage and enriching sticker diversity.
+STICKERCONV consists of 12,931 dialogue sessions, 67,505 stickers (unique 5.8K stickers) and 2K user personalities.
+Each session averages 5.22 stickers and 5.49 dialogue turns.
+Emotional label distribution analysis, as illustrated in Figure 4, highlights the differences in sticker usage between User and System, reflecting their unique roles.
+Further analysis of the STICKERCONV is detailed in the Appendix A.
 4 PEGS
-We devise an empathetic large multimodal dialog model, PEGS, with the ability to PErceive and Generate Stickers.
-Figure 4 illustrates the overall structure of our model.
-Technically, we utilize ViT-G/14 from EVA-CLIP (Fang et al., 2023), Q-Former from BLIP-2 $\text{OPT}_{\text{2.7B}}$ (Li et al., 2023b), and a linear layer to encode images.
+We devise a multimodal empathetic response generation framework, PEGS, with the ability to PErceive and Generate Stickers.
+Figure 3 illustrates the architecture of our framework.
+With different strategies to generate images, we derive three models based on this framework: PEGS-Ret/Gen/RAG, which denote the retrieval, generation, and retrieval-augmented generation methods to provide images.
+Technically, we utilize ViT-G/14 from EVA-CLIP (Fang et al., 2023), Q-Former from BLIP-2 (Li et al., 2023b), and a linear layer to encode images.
 Vicuna (Chiang et al., 2023), a widely used language model in LMMs, is employed for language modeling.
-We adopt Stable Diffusion (SD) (Rombach et al., 2022) as our image decoder.
+Stable Diffusion (SD) (Rombach et al., 2022) as the image decoder.
 4.1 Multimodal Input Perception
-Multimodal Encoding
 With reference to existing works (Li et al., 2023b; Dai et al., 2023; Zhu et al., 2023), we convert the multimodal inputs into feature vectors that can be solved by LLM.
-Specially, each text token is embedded into a vector $e_{t}\in\mathbb{R}^{1\times d}$, while each image is first encoded by a pre-trained vision encoder, and then a aligned feature vector $e_{v}\in\mathbb{R}^{32\times d}$ is obtained via Q-Former and a linear projection layer.
-Training
-Our effort is built on BLIP-2 (Li et al., 2023b), which has captured vision-language knowledge from numerous aligned image-text pairs.
-Zhu et al.
-(2023) have proven the effectiveness that freezing the pre-trained vision encoder, Q-Former, and LLM, while only pre-training a linear projection layer.
-As shown in Figure 4(a), we train the input linear projection layer using LAION115M, a sub-dataset of LAION400M (Schuhmann et al., 2021).
-The pre-training process undergoes 80,000 steps with a batch size of 32.
+Specially, each text token is embedded into a vector $e_{t}\in\mathbb{R}^{1\times d}$, while each image is first encoded by a pre-trained vision encoder, and then an aligned feature vector $e_{v}\in\mathbb{R}^{32\times d}$ is obtained via Q-Former and a linear projection layer.
 4.2 Multimodal Output Generation
 Expanding Vocabulary
-Given original LLM vocabulary $V$ of size $|V|$, to enable the LLM to determine the timing of generating visual responses and to obtain feature vectors for image generation, we extend the vocabulary with an additional tokens set $V_{\mathrm{img}}=\{\mathrm{[IMG1]},\mathrm{[IMG2]},\ldots,\mathrm{[IMG\{}r\mathrm{\}]}\}$.
-Therefore, the extended vocabulary $V^{*}=V\cup V_{\mathrm{img}}$.
+We extend the vocabulary $V$ with an additional visual tokens set $V_{\mathrm{img}}=\{\mathrm{[IMG1]},\mathrm{[IMG2]},\ldots,\mathrm{[IMG\{}r\mathrm{\}]}\}$.
 We denote the original word embedding matrix as $E\in\mathbb{R}^{|V|\times d}$.
-For the embeddings matrix $E^{*}\in\mathbb{R}^{|V^{*}|\times d}$ of the extended vocabulary $V^{*}$, the embeddings $E_{\mathrm{img}}\in\mathbb{R}^{r\times d}$ of the added special token $V_{\mathrm{img}}$ are randomly initialized and the embeddings $E$ of the original token are preserved:
-$\centering E^{*}[0:|V|,:]=E\@add@centering$
+For the embeddings matrix $E^{*}\in\mathbb{R}^{|V^{*}|\times d}$ of the extended vocabulary $V^{*}=V\cup V_{\mathrm{img}}$, the embeddings $E_{\mathrm{img}}\in\mathbb{R}^{r\times d}$ of the added special tokens are randomly initialized and the embeddings $E$ of the original token are preserved:
+$\centering E^{*}[0:|V|,:]=E\@add@centering$ (1)
+We split visual tokens into two sets, where the front $k$ tokens for image retrieval and the back $r-k$ tokens for image generation:
+$V_{\mathrm{ret}}=\{\mathrm{[IMG1]},\ldots,\mathrm{[IMG\{}k\mathrm{\}]}\}$ (2)
+$V_{\mathrm{gen}}=\{\mathrm{[IMG\{}k+1\mathrm{\}]},\ldots,\mathrm{[IMG\{}r\mathrm{\}]}\}$ (3)
+where $V_{\mathrm{ret}}$ is used in PEGS-Ret and PEGS-RAG and $V_{\mathrm{gen}}$ is used in PEGS-Gen and PEGS-RAG.
 Text Generation
@@
 The loss function $\mathcal{L}_{\mathrm{lm}}$ is defined as:
-$\centering\mathcal{L}_{\mathrm{lm}}=-\sum_{i=1}^{k}\log p(u_{i}|s,u_{1},\ldots,u_{i-1};\theta,E_{\mathrm{img}})\@add@centering$
+$\centering\mathcal{L}_{\mathrm{lm}}\!=\!-\!\sum_{i=1}^{k}\!\log p(u_{i}|s,u_{1},\ldots,u_{i-1};\theta,E_{\mathrm{img}})\@add@centering$ (4)
 where $s=\{e_{m}^{(1)},e_{m}^{(2)},\ldots,e_{m}^{(l)}\}$ and $m\in\{t,v\}$ denoting the modality.The original LLM weights $\theta$ are kept frozen, and we only update $E_{\mathrm{img}}$.
+Image Retrieval
+For image retrieval, PEGS aligns the hidden states $h_{\mathrm{ret}}$ corresponding to $V_{\mathrm{ret}}$ into the retrieval space by contrastive learning (Chopra et al., 2005).
+${W}_{\mathrm{t}}\in\mathbb{R}^{d\times e}$ and ${W}_{\mathrm{i}}\in\mathbb{R}^{p\times e}$to bridge the semantic gap and adjust the dimension.
+Cosine similarity is used to measure the similarity of the projection vectors:
+$\centering\text{sim}(x,y)=\frac{({W}_{\mathrm{t}}^{T}h_{\mathrm{ret}}(x))^{T}({W}_{\mathrm{i}}^{T}\nu_{\phi}(y))}{\|{W}_{\mathrm{t}}^{T}h_{\mathrm{ret}}(x)\|\|{W}_{\mathrm{i}}^{T}\nu_{\phi}(y))\|}\@add@centering$ (5)
+where $\nu_{\phi}$ is the image encoder.
+The projection vectors are used to minimize the InfoNCE loss (Oord et al., 2018), which consists of text-to-image (t2i) and image-to-text (i2t) loss in a batch of N text-image pairs ($x_{i}$, $y_{i}$):
+$\mathcal{L}_{\mathrm{t2i}}\!$ $=\!-\frac{1}{N}\!\sum_{i=1}^{N}\!\left(\!\log\!\frac{\exp(\text{sim}(x_{i},y_{i})\!/\!\tau)}{\sum_{j=1}^{N}\!\exp(\text{sim}(x_{i},y_{j})\!/\!\tau)}\!\right)$ (6)
+$\mathcal{L}_{\mathrm{i2t}}\!$ $=\!-\frac{1}{N}\!\sum_{i=1}^{N}\!\left(\!\log\!\frac{\exp(\text{sim}(x_{i},y_{i})\!/\!\tau)}{\sum_{j=1}^{N}\!\exp(\text{sim}(x_{j},y_{i})\!/\!\tau)}\!\right)$ (7)
+$\centering\mathcal{L}_{\mathrm{ret}}=\frac{1}{2}(\mathcal{L}_{\mathrm{t2i}}+\mathcal{L}_{\mathrm{i2t}})\@add@centering$ (8)
+$\mathcal{L}_{\mathrm{ret}}$ is the loss used to optimize the projection layers for retrieval.
 Image Generation
-Inspired by GILL (Koh et al., 2023), we align the hidden states $h_{\mathrm{gen}}$ corresponding to the output visual tokens $\{\mathrm{[IMG\{}i{\}]}\}_{i=1}^{r}$ into the input space of the image decoder.
+We align the hidden states $h_{\mathrm{gen}}$ corresponding to the output visual tokens $V_{\mathrm{gen}}$ into the input space of the image decoder.
 Specifically, we connect them through a feature mapper module, containing two linear layers and a 4-layer encoder-decoder transformer model with a learnable queries feature $q$.
-For the given image caption $c$ (and its emotion $e$ if available), our target is to minimize the MSE loss between their embeddings derived from the frozen pre-trained SD text encoder $\tau$ and the projected representations:
-$\centering\mathcal{L}_{\mathrm{gen}}=\lVert\theta_{\mathrm{Mapper}}(h_{\mathrm{gen}},q)-\tau(c,[e])\rVert_{2}^{2}\@add@centering$
-Training
-Distinguished from realistic images, stickers have abstract nature and rich emotional expressions.
-Generic text-to-image models perform well in various scenarios, but they are flawed in sticker generation.
-Constrained by limited sticker data, there are not enough image-text pairs to turn the base text-to-image model to the style of stickers, and to pre-train for multimodal generation.
-As a remedy, we adopt a variant of CuteYukiMix [^1] https://civitai.com/models/28169/cuteyukimixadorable-style, a SD model trained on cute cartoon data, as the pre-trained text-to-image model.
-In the pre-training phase, our objective is to align the representations with the text embeddings (Figure 4(b)), so we train our model on the captions from LAION115M.
-The pre-training process undergoes 50,000 steps with a batch size of 32.
-However, pre-trained SDs (focusing on obvious features, such as characters’ bodies, costumes) are insensitive to sentiment conditions, and it is difficult to directly perform end-to-end instruction tuning to learn emotion elements, we inserted a supervise fine-tuning (SFT) procedure on the sticker-text pairs with emotion to efficiently cause them to focus on emotion elements.
+For the given image caption $c$ (and its emotion $e$ if available), our target is to minimize the MSE loss between their embeddings derived from the frozen pre-trained SD text encoder $\eta$ and the projected representations:
+$\centering\mathcal{L}_{\mathrm{gen}}=\lVert\theta_{\mathrm{Mapper}}(h_{\mathrm{gen}},q)-\eta(c,[e])\rVert_{2}^{2}\@add@centering$ (9)
+Retrieval-Augmented Image Generation
+Intuitively, continuing to do generation on retrieved images can extend the diversity of images while maintaining image quality, thus we explore retrieval-augmented generation.
+Specifically, we retrieve an image serving as a latent representation $c_{I}$ for augmenting the generation process. During image generation, $h_{\mathrm{gen}}$ remains utilized as a condition.
 4.3 Joint Learning
-Pre-trained LLMs have powerful text interaction capability, but perform poorly on open-domain conversations.
-We further utilize the constructed StickerConv for jointly fine-tuning of the entire model to achieve the capabilities of open-domain empathetic multimodal perception and generation.
-The instruction prompts we used are shown in Appendix C.4.
-As shown in Figure 4(c), we train our model in an end-to-end manner, using LoRA (Hu et al., 2021) to synchronize the update of a limited number of parameters in the LLM with the input linear projection layer and the feature mapper.
+Pre-trained LLMs excel in text interactions yet struggle in empathetic conversations.
+We further utilize the constructed STICKERCONV for jointly fine-tuning of the entire model to achieve the capabilities of multimodal perception and generation.
+We train our model in an end-to-end manner, using LoRA (Hu et al., 2021) to synchronize the update of a limited number of parameters in the LLM with the input linear projection layers and the feature mappers.
 The overall loss function $\mathcal{L}$ can be represented as:
-$\centering\mathcal{L}=\lambda_{1}\times\mathcal{L}_{\mathrm{lm}}+\lambda_{2}\times\mathcal{L}_{\mathrm{gen}}\@add@centering$
-where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters. For $\mathcal{L}_{\mathrm{gen}}$, one of its inputs — emotion $e$ is available.
+$\centering\mathcal{L}=\lambda_{1}\times\mathcal{L}_{\mathrm{lm}}+\lambda_{2}\times\mathcal{L}_{\mathrm{gen}}+\lambda_{3}\times\mathcal{L}_{\mathrm{ret}}\@add@centering$ (10)
+where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ represent hyperparameters.
+$\mathcal{L}_{\mathrm{gen}}$ includes emotion $e$ as an input.
+For PEGS-Ret, $\mathcal{L}_{\mathrm{gen}}=0$, and for PEGS-Gen, $\mathcal{L}_{\mathrm{ret}}=0$.
+Implementation details can be found in Appendix C.
+5 Evaluation Metrics
+5.1 Text Metrics
+To comprehensively evaluate the fluency, diversity and accuracy of dialogue generation, we utilize an array of broadly recognized text metrics, comprising BLEU Papineni et al.
+(2002), Distinct-n (Dist-n) Li et al.
+(2016) , ROUGE-L (ROU_L) Kingma and Ba (2017), METEOR (MET) Banerjee and Lavie (2005), CIDEr Vedantam et al.
+(2015), BERTScore (BERTS) Zhang et al.
+(2020).
+5.2 Multimodal Metrics
+We use MM-Relevance (MMr) Feng et al.
+(2023) to assess the relevance between the predicted multimodal response and the golden response.
+However, MMr overlooks the frequency of multimodal replies, a critical aspect considering that text responses within a modality often show higher similarity than across modalities.
+This could result in models that rarely engage in multimodal responses receiving inaccurately high MMr scores.
+To address this issue, we impose penalties on models that are more inclined to generate text-only replies:
+$f\text{-MMr.}=(1-\alpha(1-f))\cdot\text{MMr.}$. (11)
+where $\alpha$ represents the penalty coefficient. In our experiments, we set $\alpha$ to 0.8. $f$ denotes the relative frequency at which the model produces multimodal responses, $f\in[0,1]$.
+Model BLEU-1/2/3/4 Dist-1/2/3 ROU_L. MET. CIDEr BERTS.
+Vicuna-text 0.44/0.30/0.22/0.17 0.879/0.994/0.999 0.31 0.37 0.39 0.878
+Vicuna-tool 0.43/0.29/0.22/0.17 0.870/0.989/0.994 0.30 0.36 0.38 0.900
+ChatGLM3-text 0.42/0.28/0.21/0.16 0.806/0.981/0.996 0.31 0.40 0.40 0.886
+ChatGLM3-tool 0.36/0.22/0.16/0.11 0.859/0.992/0.998 0.26 0.34 0.20 0.899
+PEGS-Ret 0.46/0.32/0.25/0.20 0.839/0.989/0.997 0.34 0.42 0.47 0.906
+PEGS-RAG 0.46/0.32/0.25/0.20 0.839/0.989/0.997 0.34 0.42 0.47 0.906
+PEGS-Gen 0.47/0.33/0.26/0.21 0.848/0.990/0.997 0.35 0.44 0.57 0.911
+Table 1: Results of quality of text generate in PEGS and baseline models.
+5.3 LLM-based Metrics
+LLMs are capable of grading similarly to humans, providing scores for both textual and sticker outputs, thus enabling a comprehensive multimodal evaluation system.
+We introduce three LLM-based metrics: (1) Empathy: We assess empathy in model responses, both textual (Empathy-text, EMP-txt) and multimodal (Empathy-multimodal, EMP-mm), averaging scores from five independent scorings to reduce randomness and bias.
+(2) Consistency: Based on the context, we assign consistency scores for textual and sticker responses, marked as Consistency (CON), employing the same scoring method.
+(3) Rank: We compare responses of different models to the same context.
+The ranker organizes responses based on quality, empathy, and consistency, averaging across many possible ranking combinations to ensure fairness and objectivity.
+5.4 Human Metrics
+Considering the subjectivity of empathy and the complexity of multimodal responses, we establish seven detailed manual evaluation metrics: Sticker Generation Quality (StiGQ), Empathy Sticker (Es), Empathy Text (Et), Consistency (Con), Fluency (Flu), Informativity (Inf), Sticker Diversity (StiD).
+Metrics are scored on a 1 to 5 scale, with higher scores denoting better performance.
+For fairness and consistency, all models are evaluated in identical contexts.
+The evaluation panel consists of five members who score models anonymously, unaware if they are assessing a baseline model.
+The human evaluation process randomly selects 100 dialogue sessions from the STICKERCONV test set.
 6 Experiments
 Experiments are conducted on STICKERCONV. We perform response predictions for all turns of each dialogue and consider all previous turns as context.
-Models IP EX ER diff-IP diff-EX diff-ER dist-1 dist-2 dist-3
-Vicuna-7B 0.1857 0.2405 0.9509 0.3681 0.9745 0.2651 0.7995 0.9783 0.9936
-ChatGLM3-6B 0.1632 0.2699 0.9768 0.344 0.9993 0.2629 0.7854 0.9776 0.9939
-PEGS 0.1723 0.4458 0.9724 0.3113 0.8721 0.244 0.8397 0.9888 0.9976
-Table 1: Results of quality of text generate in PEGS and baseline.
-5.1 Experimental Settings
-Implementation Details
-At each training stage, the AdamW optimizer is adopted, and the cosine annealing scheduler is used to adjust the learning rate with the initial learning rate of 1e-4.
-The details of pre-training have been introduced in Section 4.
-Our model is fine-tuned on the entire StickerConv.
-It is noteworthy that only 3% of the model parameters are updated, and it takes about 6 hours on a single A800 GPU.
-Baseline
-We compare with the following baselines: (1) Vicuna-7B: An open-source chatbot trained by fine-tuning LLaMA on user-shared conversations collected from ShareGPT Chiang et al.
+6.1 Implementation Details
+Our models are pre-trained on LAION115M, a sub-dataset of LAION400M (Schuhmann et al., 2021), and fine-tuned on STICKERCONV.
+More implementation details and hyperparameter settings are provided in Appendix C.
+The training procedure is conducted on 2 NVIDIA A6000 48G GPUs.
+6.2 Baselines
+We benchmark against Vicuna-7B Chiang et al. (2023) and ChatGLM3-6B Zeng et al. (2023) models, employing two experimental paradigms:
+Text Fine-tuning
+We fine-tune the models using text data from the STICKERCONV training set and then make predictions on the test dataset.
+Tool Learning
+We fine-tune the models with multimodal data from the STICKERCONV training set.
+We use stable-diffusion as a tool to teach the models how to utilize it for generating sticker responses Tang et al.
 (2023).
-We use the Vicuna-delta-v0 version to ensure consistency with the model base used during training.
-(2) ChatGLM3-6B: An open-source, bilingual supported large-scale dialogue model Zeng et al.
-(2023).
-5.2 Evaluation Metrics
-5.2.1 General Metrics
-This category focuses on assessing the quality of the model’s textual output.
-Diversity
-We adopt Distinct-n (Dist-1, 2, 3) Li et al. (2016) as quantitative indicators of diversity, measuring the variety and creativity of the model’s outputs.
-Empathy Performance
-To evaluate the model’s capacity for empathy, we adopt the EPITOME Sharma et al. (2020a) method, comprising three sub-metrics: IP (interpretations), EX (explorations) and ER (emotional reactions).
-Empathy Difference
-DIFF-EPITOME Lee et al. (2022) measures the difference scores of IP, EX, ER between the golden response and predicted response.
-5.2.2 LLM-based Metrics
-LLM has been widely proven to be capable of acting as a grader similar to humans and can provide reasonable scores based on the designed prompts.
-These metrics not only focues on textual outputs but also include sticker outputs, forming a comprehensive multimodal evaluation system.
-Empathy
-Utilizing ChatGPT-3.5-turbo as a scorer, based on the context, we assess empathy scores for the model’s textual and sticker responses, denoted as Emp-txt and Emp-sticker, respectively.
-To mitigate randomness and bias, we conduct five independent scorings for each response and calculate the average.
-Consistency
-Again using ChatGPT-3.5-turbo as a scorer, based on the context, we assign consistency scores for textual and sticker responses, marked as Con-txt and Con-sticker, employing the same scoring method.
-Ranking
-Employing ChatGPT-3.5-turbo as a result ranker, we compare responses of different models to the same context.
-The ranker organizes responses based on quality, empathy, and consistency, averaging across all possible ranking combinations to ensure fairness and objectivity in assessment.
-5.2.3 Multimodal Metrics
-CLIP-Score
-We utilize CLIP (Radford et al., 2021) as multimodal encoder, evaluate the consistency of model’s multimodal responses.
-MM-Relevance
-MM-Relevance Feng et al.
-(2023) is a metric designed to measure the relevance of multimodal responses.
-Traditional mm-relevance metrics, however, face limitations, particularly in cases where there is a lack of alignment between model responses and golden responses.
-To address these challenges, we propose Inter-MM-Rel (Interactive multimodal Relevance).
-This innovative approach calculates and averages the similarity scores between each mode of information, without the constraint of alignment, allowing for a more comprehensive and accurate assessment of multimodal response relevance.
+To ensure fairness, the tool models employ the same SD model as PEGS.
 6.3 Result Analysis
-From Table 1, PEGS shows remarkable performance in several metrics, particularly in EX (Explorations), DIFF-EPITOME and diversity measuresm indicating a superior ability in generating diverse and exploratory responses.
-However, PEGS’s lower scores in IP, ER could be due to two reasons: (1) The general metrics only evaluate the text output and underestimate the actual empathy ability of the PEGS.
-(2) In multimodal empathy dialogue task, stickers play a significant role in conveying emotional reactions and explanations, while text handles more explorative tasks.
-Models Rank Emp-txt Emp-sticker Con-txt Con-sticker
-Vicuna 2.12 3.6768 4.3221
-ChatGLM3 1.99 3.6908 4.3405
-PEGS 1.88 3.7354 4.2534 4.2739 4.1155
-Table 2: Results of LLM-Scorer.
-Table 2 reveals PEGS’s strong performance in ranking, empathy text (Emp-txt), empathy sticker, and consistency of sticker responses (Con-sticker).
-However, its lower consistency in text (Con-txt) might be attributed to its emphasis on empathetic stickers, which could compromise some textual context consistency.
-This is in line with the additional information suggesting that the empathetic effectiveness of stickers might require some trade-offs in textual context consistency.
-Models MM-Rel Inter-MM-Rel CLIP-Score
-PEGS 69.41 59.06 21.03
-Table 3: Results of multimodal evaluate.
-Table 3 showcases the multimodal capabilities of PEGS.
-PEGS performs well in multimodal relevance (MM-Rel and Inter-MM-Rel) and consistency (CLIP-Score).
-This suggests that PEGS effectively integrates different modalities in its responses, maintaining relevance and consistency across them.
+Table 1 reports the results for text metrics.
+PEGS’s performance on the Dist-n is slightly lower than that of the Vicuna.
+This may be attributed to PEGS expanding Vicuna’s vocabulary to facilitate sticker perception and generation, impacting PEGS’s text diversity to a certain degree.
+In our experiment, the training objective of tool learning includes both text generation and tool calling.
+These two objectives may interfere with each other.
+As a result, the tool model’s text capabilities fall short of the text fine-tuning model’s.
+PEGS’s end-to-end structure integrates multimodal inputs and outputs, streamlines tasks with unified training objectives, and yields the best text results.
+PEGS-Gen utilizes fewer special tokens (32), thereby outperforming PEGS-Ret and PEGS-RAG in text metrics.
+These findings corroborate the efficacy of PEGS framework in generating text responses of high quality and accuracy.
+Model Freq. MMr. f-MMr.
+Vicuna-tool 0.141 0.725 0.602
+ChatGLM3-tool 0.905 0.659 0.647
+PEGS-Ret 0.850 0.674 0.653
+PEGS-RAG 0.847 0.680 0.659
+PEGS-Gen 0.811 0.672 0.647
+Table 2: Results of multimodal metrics.
+Table 2 displays the results for multimodal metrics, where Freq represents the relative frequency of stickers replies from each model.
+Although Vicuna-tool achieves a high MMr (72.48), this result is largely due to its lower Freq (0.141), which may not accurately represent its multimodal interaction capabilities.
+F-MMr is more robust than MMr because it accounts for Freq.
+PEGS excels in f-MMr, showcasing that its end-to-end structure, integrating text and stickers, achieves high consistency in multimodal reply generation.
+Model EMP-txt EMP-mm CON. Rank
+Vicuna-text 3.677 - 4.322 4.527
+Vicuna-tool 3.822 3.799 4.247 4.093
+ChatGLM3-text 3.691 - 4.341 5.037
+ChatGLM3-tool 3.700 3.760 4.220 4.400
+PEGS-Ret 3.873 4.040 4.380 4.030
+PEGS-RAG 3.927 4.076 4.370 3.900
+PEGS-Gen 3.768 4.353 4.404 1.917
+Table 3: Results of LLM-based metrics.
+Table 3 presents the evaluation results for the LLM.
+Relative to the baseline model, the two tool-learning models outperform the text model in ranking.
+This underscores the pivotal role of stickers in enhancing empathetic communication.
+According to the results from LLM-based metrics, PEGS can produce empathetic texts while ensuring high consistency, and utilize stickers to enhance emotional resonance.
+Table 2 and PEGS-Gen’s EMP-mm (4.353) indicates that the quality of the multimodal response directly correlates with the extent to which it enhances empathy.
+Model StiGQ. Et. Es. Con. Flu. Inf. StiD.
+Vicuna-tool 4.09 4.07 3.78 4.08 4.23 4.08 3.20
+ChatGLM3-tool 4.32 4.06 3.24 4.11 4.58 3.99 3.10
+PEGS-Ret - 4.11 4.17 4.22 4.36 4.09 3.40
+PEGS-RAG 3.37 4.12 2.92 4.22 4.36 4.10 3.80
+PEGS-Gen 4.53 4.29 4.19 4.37 4.47 4.26 3.60
+Table 4: Results of the human evaluation.
+Table 4 shows the result of human evaluation.
+PEGS notably outstrips the two baseline models across most evaluative indicators, save for Flu, thus corroborating the efficacy of the PEGS framework.
+The Flu of PEGS ($<$4.5) is lower compared to ChatGLM3-tool (4.58), likely because its high Inf ($>$4.09) results in a slight reduction in Flu.
+According to Inf and Et, the amount of information is a crucial factor influencing empathy.
+PEGS-RAG excels in sticker diversity (3.8) but records the lowest score in sticker generation quality (3.37).
+This observation suggests that although the RAG strategy enhances sticker variety, it might also reduce the quality of sticker generation in this experiment.
+Analysis of StiGQ, Es, and StiD reveals that multimodal responses can enhance empathy, and this enhancement is positively correlated with the responses’ quality.
+When considering all evaluation of human metrics, PEGS-Gen stands out remarkably, partly because of its minimal use of special tokens and partly due to its proficient sticker generation capability (owe to end-to-end structure).
+These outcomes highlight PEGS-Gen’s exceptional ability to produce high-quality, emotionally resonant, and diverse multimodal responses.
+Further analysis is presented in Appendix D. We collect the case study in Appendix E.
 7 Conclusion
@@
 We developed the Agent for STICKERCONV, a sophisticated LLM-based multi-agent system capable of simulating human-like interactions using stickers, thereby creating multimodal empathetic responses.
-Building upon the StickerConv, we developed PErceive and Generate Stickers (PEGS), an advanced multimodal dialogue model.
-This model is adept at both perceiving and generating stickers, effectively enhancing the communication experience.
+Building upon the STICKERCONV, we developed PErceive and Generate Stickers (PEGS), an advanced multimodal empathetic dialogue framework.
+This framework adeptly perceives and generates stickers, effectively enhancing the communication experience.
 Furthermore, we established a comprehensive set of evaluation metrics for multimodal empathy tasks based on LLM.
 We are confident this work will be a valuable asset in advancing research in the field of multimodal empathetic dialogue systems.
-Appendix A Appendix
-A.1 Dataset Statistics
+Limitations
+Although our study has achieved certain advancements, it is not without its limitations.
+Firstly, STICKERCONV is exclusively in English, implying that our model and evaluations might not be directly translatable to other linguistic contexts.
+Secondly, although we investigate the retrieval-augmented generation (RAG) variant of the PEGS, hoping to enhance the quality and diversity of sticker generation, the RAG version has demonstrated instability in generating quality stickers in practice.
+Despite these issues, the pursuit of diversity and quality in sticker generation means that the explorations into the RAG version remain valuable.
+It suggests a promising research trajectory: how to elevate the stability and variety of the generated content by refining algorithms or amalgamating additional technological solutions.
+Finally, from our experiments, PEGS and two baseline tool-model exhibit abnormal frequencies in the usage of stickers.
+Aligning the model’s sticker usage frequency and timing with human-like interaction patterns is an important area for future exploration.
+Ethics Statement
+The STICKERCONV dataset, produced by the large language model (LLM), therefore sidesteps privacy concerns.
+Nonetheless, LLMs may inherently harbor biases, and both STICKERCONV and PEGS could inadvertently magnify these biases in their application.
+Despite our efforts to mitigate these biases by incorporating a diverse range of personality information into Agent4SC, it is important to acknowledge that biases intrinsic to LLMs cannot be completely eradicated.
+This inherent bias is an ongoing challenge that needs to be addressed continuously through iterative improvement and monitoring of the model.
+PEGS reveals its robust capability in creating multimodal empathetic replies, adept at grasping human emotions and generating pertinent multimodal content.
+Such abilities, if exploited, could potentially manipulate user emotions.
+To forestall these risks, we have implemented a mechanism for detecting NSFW content in the generation of multimodal content.
+It is imperative to underscore that, despite our model’s proficiency in aiding the understanding and response to human emotions, it is not intended to entirely supplant human emotional communication.
+Excessive dependency on AI for empathetic interactions might erode human empathic abilities and the cultivation of emotional intelligence.
+Appendix A Dataset Statistics
+In our dataset creation, it is important to note that vector databases numbered 1 to 80 are exclusively employed for the creation of training and cross-validation sets, whereas those numbered 81 to 100 are reserved for the construction of the test set.
+We configure the agent’s "select" action’s top-K parameter to 10.
 Table 5 offers a statistical breakdown of the dataset, segmented into training, validation, and test splits.
 It provides a quantitative overview, detailing the number of instances in each subset, the unique stickers, turn counts, and token counts.
+Our dataset is comparable in size to PhotoChat Zang et al.
+(2021).
+To our knowledge, it is the first multimodal empathetic dialogue dataset, uniquely integrating stickers as non-textual elements to enhance empathetic communication more effectively.
 split Number Unique Sticker Turn $\text{Token}_{\text{vicuna}}$ $\text{Token}_{\text{GPT}}$
@@
 Table 5: The statistics of STICKERCONV.
+Figure 4 shows the emotional distribution of stickers used by the User and the System.
+It reveals a striking trend: users have a significant preference for stickers that convey negative emotion, in contrast to the system.
+The system predominantly utilizes stickers to express neutral and positive emotion.
+This comparison not only reflects the distinct emotional expression preferences between users and the system but also highlights the system’s active and supportive role in interactions.
+Figure 10 and 11 present two examples from STICKERCONV.
+Figure 4: The chart of emotional distribution in the choice of stickers between the User and the System.
 Figure 5 showcases the emotion distribution within user profiles in the Agent for STICKERCONV.
 The chart illustrates the percentage share of each emotion, offering insight into the prevalence of affective expressions within the dataset.
+Increasing personality distributions enhances the Agent4SC’s diversity, thereby helping to mitigate the inherent bias issue in LLMs.
+Figure 9 shows a profile example across various emotions.
 Figure 5: Emotion distribution of user profile in Agent for STICKERCONV.
@@
 Figure 6: The 200 most popular emotion-related words in STICKERCONV.
-Appendix B Case Study
-We conduct a series of case studies to provide a more straightforward illustration of the performance of our model instruction-tuned by the StickerConv in multimodal empatheric conversations.
-These cases demonstrate the ability of PEGS to empathize with users and perceive the stickers provided by users, along with generating multimodal responses with vivid stickers.
-As shown in Figure 7, in the positive emotion condition, PEGS will reward users with praise and support.
-As shown in Figure 8, two conversations in a negative scenario are displayed.
-PEGS as a conciliator can comfort users and guide their emotions towards positivity.
-Appendix C Prompts
+Appendix B Implementation Details of Agent4SC
+B.1 Tool Module
+Before advancing to the follwing process, we altered the background of the SER30K stickers from black to white. The detailed process of building the Tool module is as follows:
+Filter
+Leveraging the LLaVa-v1.5-13b (Liu et al., 2023a), we craft a nuanced Chain of Thought (CoT) Wei et al.
+(2023) procedure to meticulously scrutinize each SER30K sticker.
+This approach entails the model initially summarizing the sticker’s content, subsequently assessing its relevance to casual conversation, and finally determining suitability through iterative responses.
+Each sticker undergoes multiple evaluations to ensure accuracy.
+Annotation
+Filtered stickers undergo a second CoT analysis to answer 5 questions about their content, emotions, usage contexts, tone, and any comedic or satirical elements.
+This detailed annotation process provides deep insights, structured as Q&A pairs for further use.
+Knowledge
+Leveraging ChatGPT-3.5-Turbo, we extract critical information from the Q&A pairs, generating a comprehensive understanding of each sticker, including its description, emotion, and usage recommendation.
+Split
+We segment the filtered dataset into smaller vector databases, incorporating the original emotional labels alongside freshly acquired knowledge.
+Each database maintains an emotional sticker distribution that aligns with that of the original dataset.
+Following the process described, we obtain a sticker database containing 10,648 stickers.
+The metadata for each sticker includes the original emotion label, the description of emotion, the description of content, and usage recommendations.
+The sticker database is then evenly distributed into 100 smaller vector databases based on emotional balance.
+B.2 Manager Agent
+During dataset creation, we limit conversation turns to a maximum of six.
+The manager agent is programmed to review the conversation every two turns, and to verify the consistency of each sticker used with the conversation context.
+We observe that with more conversation turns, the likelihood of the agent repeating previous statements increased.
+Upon detecting such repetition, the manager agent will transition the agent’s LLM to GPT-4 to prevent further repetition.
+Model learning rate warmup steps weight decay batch size max length steps
+Input PEGS 1e-4 2,000 0.05 128 32 80,000
+Output PEGS-Ret 3e-5 4,000 1e-4 36 77 142,220
+PEGS-RAG 3e-5 4,000 1e-4 36 77 142,220
+PEGS-Gen 1e-5 1,000 0.05 108 32 80,000
+Table 6: Hyperparameters for pre-training PEGS.
+Model learning rate warmup steps weight decay batch size max length epochs
+PEGS-Ret 3e-5 70 0.05 36 77 10
+PEGS-RAG 3e-5 70 0.05 36 77 10
+PEGS-Gen 5e-5 200 0.05 108 32 4
+Table 7: Hyperparameters for fine-tuning the output side of the PEGS on sticker-text pairs.
+Model learning rate warmup steps weight decay LoRA-r/alpha/dropout batch size max length epochs
+Vicuna-text 5e-4 300 0.00 8/16/0.05 16 4,096 5
+Vicuna-tool 5e-4 300 0.00 8/16/0.05 16 4,096 5
+ChatGLM3-text 1e-4 300 0.10 8/16/0.05 16 4,096 5
+ChatGLM3-tool 1e-4 300 0.10 8/16/0.05 16 4,096 5
+PEGS-Ret 3e-6 2,000 0.05 8/16/0.05 4 768 9
+PEGS-RAG 3e-6 2,000 0.05 8/16/0.05 4 768 9
+PEGS-Gen 1e-4 1,000 0.05 8/16/0.05 4 768 4
+Table 8: Hyperparameters for fine-tuning baselines and PEGS on STICKERCONV.
+Appendix C More Implementation Details of PEGS
+C.1 Pre-training
+Multimodal Perception
+Our effort is built on BLIP-2 (Li et al., 2023b), which has captured vision-language knowledge from numerous aligned image-text pairs.
+Zhu et al.
+(2023) have proven the effectiveness that freezing the pre-trained vision encoder, Q-Former, and LLM, while only pre-training a linear projection layer.
+We train the input linear projection layer using LAION115M.
+Multimodal Generation
+Distinguished from realistic images, stickers have an abstract nature and rich emotional expressions.
+Generic text-to-image models perform well in various scenarios, but they are flawed in sticker generation.
+Constrained by limited sticker data, there are not enough image-text pairs to turn the base text-to-image model to the style of stickers, and to pre-train for multimodal generation.
+As a remedy, we adopt a variant of CuteYukiMix [^2] https://civitai.com/models/28169?modelVersionId=163923, a SD model trained on cute cartoon data, as the pre-trained text-to-image model.
+In the pre-training phase, our objective is to align the representations with the text embeddings, so we train our model on the captions from LAION115M.
+The AdamW optimizer is adopted, and the cosine annealing scheduler is used to adjust the learning rate. Table 6 shows the hyperparameter settings in the pre-training phase.
+C.2 Fine-tuning
+Multimodal Generation
+Pre-trained diffusion models (focusing on obvious features, such as characters’ bodies, costumes) are insensitive to sentiment conditions, and it is difficult to directly perform end-to-end fine-tuning to learn emotion elements, we inserted a supervise fine-tuning (SFT) procedure on the image-text pairs with emotion to efficiently cause them to focus on emotion elements.
+Table 7 shows the hyperparameter settings for this phase.
+Joint Learning
+We utilize the constructed STICKERCONV for fine-tuning of the entire model to achieve the capabilities of multimodal perception and generation.
+We train our models in an end-to-end manner.
+The AdamW optimizer is adopted, and the cosine annealing scheduler is used to adjust the learning rate.
+Table 8 shows the hyperparameter settings in the joint-learning phase.
+In this work, $\lambda_{1}=1$, $\lambda_{2}=1$, $\lambda_{3}=1$ in Eq 10.
+Appendix D Further Result Analysis
+Table 8 presents the hyperparameters for fine-tuning baselines and PEGS.
+The end-to-end architecture of PEGS necessitates significant graphics memory during training, limiting it to smaller batchsize (4) and maximum lengths (768).
+Conversely, the baseline model accommodates larger batchsize (16) and maximum lengths (4096).
+These parameters suggest that PEGS can be further optimized for use on devices with larger graphics memory.
+D.1 Analysis of Multimodal Metrics
+From Table 2, taking into account the relative frequency of multimodal responses, f-MMr proves to be more reasonable and robust compared to MMr.
+However, the multimodal response frequency of all models appears to be unusual, with the Vicuna-tool’s frequency being notably low (0.141) and that of other models excessively high ($>$0.8).
+For comparison, the System’s sticker usage frequency in the STICKERCONV is 0.6.
+This outcome suggests that future research should focus on adjusting the multimodal response frequency and timing of models to better suit human interaction.
+In this paper, we do not use traditional image generation evaluation metrics, such as FID (Heusel et al., 2017), IS (Salimans et al., 2016).
+These metrics are applied to evaluate the quality and diversity of realistic images.
+FID and IS are obtained by pre-trained Inception v3 (Szegedy et al., 2016) (pre-trained on realistic images) after certain processing.
+However, sticker possesses an abstract style and characteristics that do not match the application domains of FID and IS.
+Based on the above facts, we evaluate the quality and diversity of image generation by human evaluation.
+D.2 Analysis of LLM-based Metrics
+When using an LLM for scoring, to eliminate randomness, we repeat the scoring for each metric 5 times and calculate the average to determine the final score.
+From Table 3, the EMP-mm of PEGS outperforms the two tool models, likely owing to PEGS’s end-to-end framework that ensures greater consistency between text and stickers.
+The top-ranking performance of PEGS-Gen further demonstrates the positive impact of text-sticker relevance on empathy.
+D.3 Analysis of Human Metrics
+Figure 12 shows the complete questionnaire of human evaluation, which includes the detailed description and scoring criteria for each metric.
+The five scorers are graduate students in computer science, specializing in research on emotional reasoning or empathetic dialogue, with a thorough understanding of empathy tasks.
+During scoring, they are unaware of which model serves as the baseline and which is the PEGS.
+Metric StiGQ. Et. Es. Con. Flu. Inf. StiD.
+ICC 0.920 0.540 0.928 0.457 0.728 0.608 0.365
+Table 9: ICC Analysis Results for Human Evaluators
+Table 9 shows the results of the inter-rater reliability analysis conducted on the five scorers involved in the manual evaluation, using the Intraclass Correlation Coefficient (ICC) WEIR (2005) as the assessment criterion.
+Human scorers demonstrate high consistency in metrics like StiGQ, Es, Flu and Inf.
+The ICC of Et scores are moderate, likely because assessing empathy through text is highly subjective and depends significantly on individual life experiences.
+Similarly, the moderate ICC scores for Con may stem from the nature of our task as a multimodal empathetic response task.
+Both the context and response involve multimodal information (sticker), and the text is affected by the stickers.
+This can, to some extent, disrupt the scorers’ assessment of Con.
+The ICC of StiD is poor ($<$0.4), primarily due to it being a global metric with a single score per model, leading to higher volatility.
+Metrics EMP-txt EMP-mm CON. Rank ID. f-MMr.
+Spearman 0.40 0.70 0.87 0.40 0.80 0.90
+Kendall 0.40 0.60 0.74 0.40 0.60 0.80
+Table 10: Correlation between the metrics and human annotation.
+Table 10 illustrates the correlation between various evaluative metrics and human annotations, as evidenced by Spearman’s and Kendall’s correlation coefficients.
+The EMP-txt metric calculates the correlation with the Et of human metrics, whereas EMP-mm corresponds to the Es.
+CON is matched with the Con, and Rank correlation is determined by averaging all scores from human evaluations.
+The ID metric measures correlation with the StiD of human metrics, and f-MMr matches the average of Con, Flu, Inf, and StiD.
+Both CON and f-MMr exhibit the highest consistencies under two coefficients, achieving 0.87 and 0.90 for Spearman and 0.74 and 0.80 for Kendall, respectively.
+These results emphasize the precision and uniformity of these metrics in mirroring human assessments.
+EMP-mm and ID metrics also exhibit substantial correlations, suggesting a significant alignment with human perceptions within these domains.
+Calculating the correlation between the results of LLM-based metrics and the results of human metrics demonstrates that with carefully designed prompts, LLM can effectively evaluate empathy tasks.
+In summary, our experiments show that PEGS excels in producing varied and richly informative empathetic text responses, ensuring high contextual consistency, and generating diverse, high-quality stickers to enhance empathetic effects.
+These findings highlight the benefits of PEGS’s end-to-end architecture.
+The success of the PEGS demonstrates the effectiveness of Agent4SC and STICKERCONV.
+Considering the cost of Agent4SC (shown in Appendix G), once PEGS training is completed, inference can be deployed locally to balance cost-effectiveness and efficiency.
+This paper establishes a valuable resource for future research on multimodal empathetic response generation.
+Appendix E Case Study
+To further investigate the quality of the multimodal responses predicted by our proposed framework, we show an example of STICKERCONV test data in Figure 7.
+As we have seen, the textual responses generated by multimodal baselines and our models are consistent with the dialogue context in terms of content and emotion, demonstrating the effectiveness of fine-tuning based on STICKERCONV.
+All models provide visual responses (stickers) within the given context.
+The difference is that PEGS has more pronounced emotional tendencies, which facilitates empathy.
+To illustrate more straightforwardly the performance of our model in multimodal empathic conversations, we show several cases of user-chatbot interaction.
+As shown in Figure 8, these cases demonstrate the ability of PEGS to empathize with users and perceive the stickers provided by users, along with generating multimodal responses with vivid stickers.
+For the positive dialogue context, our model generates positive stickers to enhance interactivity and emotional expression.
+For the negative dialogue context, our model reassures the user by an appropriate sticker (the hugging behavior in Figure 8 right).
+Figure 7: An example of STICKERCONV test set.
+Top: the multimodal dialogue context between “User” and “System”.
+Bottom: the multimodal responses generated by our designed multimodal baselines and models.
+Figure 8: Examples of conversations by users interacting with PEGS.
+Users can chat with multimodal content (text and stickers) and will receive multimodal empathetic responses.
+Left: a conversation characterized by positive emotion (happiness).
+Right: a conversation characterized by negative emotion (sadness).
+Appendix F Prompts
 All {format_instructions} in the prompt utilize the response of langchain [^3] https://www.langchain.com/’s response_schema function implements format control.
 For detailed formatting information, please refer to our project’s open-source repository: https://github.com/ZhangYiqun018/StickerConv.
-C.1 Sticker Process
+F.1 Sticker Process
 Figure 13 shows the prompt of sticker process.
 For each sticker, we subject it to a process of filtering, annotation, and knowledge extraction, ultimately yielding three pieces of information: description, emotion and recommendation.
-C.2 Sticker Agent Chat
+F.2 Sticker Agent Chat
 Figure 14 shows the chat prompt, the summary encompasses the persona of the user or system, the current status, and the core traits generated based on memory.
-Figure 11, 12, 13 show the process of used sticker.
-C.3 LLM-Based Scorer
+Figure 15, 17, 18 show the process of used sticker.
+F.3 Manager Agent
+Figure 16 shows the prompt of manager agent. By using various "format instructions", this prompt can serve multiple purposes, including quality and consistency checks, among others.
+F.4 LLM-Based Scorer
 Figure 19 illustrates the template for the preference rank scorer.
@@
 It then proceeds to rank these responses and provide reasoning for the assigned rankings.
-Figures 15 and 16 depict the templates for the consistency scorer.
-The selection of the appropriate template depends on the modality of the response generated by the model.
-Specifically, Figure 15 is utilized for scoring purely textual responses.
-In contrast, when the response is multimodal, the scoring is conducted using the template shown in Figure 16.
-Similarly, Figures 17 and 18 present the templates for the empathy scorer.
-The methodology for processing responses of different modalities aligns with the approach described for the consistency scorer.
-C.4 Joint Learning
-Figure 19 shows the instruction prompts for joint learning.
-We employed an instruction template similar to Vicuna_delta_v0.
-For every query or response that uses a sticker, we append an <IMG> identifier at the end.
-Figure 7: Examples of conversations with positive emotions. Users can chat with multimodal content (text and stickers) and will receive positive multimodal responses.
-Figure 8: Examples of conversations with negative emotions.
-Our model can empathize with users who are suffering from negative emotions, such as sadness (left) or anger (right) .
-Users will be comforted and guided to positive sentiments.
+Figures 20 depict the templates for the consistency scorer. It is utilized for scoring purely textual responses.
+Figures 21 and 24 present the templates for the empathy scorer. The methodology for processing responses of different modalities aligns with the approach described for the consistency scorer.
+F.5 Joint Learning
+Figure 22 shows the prompt template for joint learning. We employed an instruction template similar to Vicuna v1.5. For every dialogue that uses a sticker, we append an <IMG> identifier.
+F.6 Tool Learning
+Figure 23 presents the instruction prompts for tool learning.
+To instruct the LLM in utilizing the SD model for sticker generation, we transform the STICKERCONV to tool format.
+When the model requires invoking the SD model, it outputs "ASSISTENT Action" along with the corresponding prompt ("ASSISTENT Action Input").
+We then feed this prompt into the pre-configured SD model and relay the outcomes back to the LLM as "ASSISTENT Observation".
+Appendix G Costing
+Extracting knowledge from labeled stickers in Llava using ChatGPT-3.5-turbo, which includes stickers in SER30K dataset, stickers generated by tool models and PEGS, costs approximately \$130.
+The construction of 2000 user profiles for Agent4SC with ChatGPT-3.5-turbo incurred a total cost of \$5.5.
+Constructing the STICKERCONV dataset with ChatGPT-3.5-turbo incurred a total cost of \$438.53, covering training, validation, and testing sets, at an average of \$0.03 per dialogue session.
+According to our test set, evaluating metrics other than rank for each comparison model costs \$22. Combining all model responses with the evaluation Rank metric costs about \$60.
+Figure 9: Example Profile of Agent4SC.
+Figure 10: Example I from STICKERCONV.
+Figure 11: Example II from STICKERCONV.
+Figure 12: Human Evaluation Questionnaire.
 Figure 13: Prompt Templates for Process Sticker.
 Figure 14: Prompt Template for Chat Prompt.
+Figure 15: Prompt Template for Manager Agent.
 Figure 16: Prompt Template for Sticker Intent.
@@
 Figure 19: Prompt Template for Ranking score.
-Figure 15: Prompt Template for Consistency Scorer (text only).
-Figure 16: Prompt Template for Consistency Scorer (multimodal).
+Figure 20: Prompt Template for Consistency Scorer.
 Figure 21: Prompt Template for Empathy Scorer (text only).
+Figure 22: Prompt Template for Joint Learning.
+Figure 23: Prompt Template for Tool Learning.
 Figure 24: Prompt Template for Empathy Scorer (multimodal).
-Figure 19: Prompt Template for Joint Learning.